### v1

# Customizing Text-to-Image Diffusion with Camera Viewpoint Control

Abstract
* indicates equal contribution
Model customization introduces new concepts to existing text-to-image models, enabling the generation of the new concept in novel contexts. However, such methods lack accurate camera view control w.r.t the object, and users must resort to prompt engineering (e.g., adding “top-view”) to achieve coarse view control. In this work, we introduce a new task – enabling explicit control of camera viewpoint for model customization. This allows us to modify object properties amongst various background scenes via text prompts, all while incorporating the target camera pose as additional control. This new task presents significant challenges in merging a 3D representation from the multi-view images of the new concept with a general, 2D text-to-image model. To bridge this gap, we propose to condition the 2D diffusion process on rendered, view-dependent features of the new object. During training, we jointly adapt the 2D diffusion modules and 3D feature predictions to reconstruct the object’s appearance and geometry while reducing overfitting to the input multi-view images. Our method outperforms existing image editing and model personalization baselines in preserving the custom object’s identity while following the input text prompt and the object’s camera pose.
Figure 1: Given multi-view images of a new object (left), we create a customized text-to-image diffusion model with camera pose control. The customized model can synthesize the object (denoted as V ¡category name¿) in novel appearances and scenes, such as A green V car, or A beetle-like V car, while allowing users to specify the target camera pose for the object. We can generate panorama images while controlling the camera pose for each object or compose multiple concepts using MultiDiffusion [3] with our model.
1 Introduction
Recently, we have witnessed an explosion of works on customizing text-to-image models [74, 23, 45, 16]. They allow us to quickly acquire visual concepts, such as personal objects and favorite places, and reimagine them with new environments and attributes. For instance, we can customize a model on our Teddy bear and prompt it with “Teddy bear on a bench in the park.” Unfortunately, customization methods lack precise camera pose control, as existing diffusion models are trained purely on 2D images without ground truth camera poses. As a result, users often rely on text prompts such as “front-facing” or “side-facing”, a tedious and unwieldy process to control views.
What if we wish to synthesize a new object, e.g., the Teddy bear in Figure 1, in a different context while controlling its pose? In this work, we introduce a new task: given multi-view images of the object, we customize a text-to-image model while enabling precise control of the new object’s camera pose. During inference, our method offers the flexibility of conditioning the generation process on both a target pose and a text prompt.
Neural rendering methods have allowed us to accurately control the 3D viewpoint of an existing scene, given multi-view images [39, 57, 5, 4]. Similarly, we seek to imagine the object from novel viewpoints but in a new scene. However, as pre-trained diffusion models, such as Latent Diffusion models [72], are built upon a purely 2D representation, connecting the 3D neural representation of the object to the 2D internal features of the diffusion model remains challenging.
In this work, we propose a new method, CustomDiffusion360, to bridge the gap between 3D neural capture and 2D text-to-image diffusion models by providing additional camera pose control w.r.t. the new custom object in 2D text-to-image models. More concretely, given multi-view images of an object, we learn to predict neural feature fields in the intermediate feature spaces of the diffusion model U-Net. To condition the generation process on a target pose, we render the features of the predicted feature fields and then fuse them with the target pose’s noisy features. This new conditioning module is added to a subset of transformer layers in the pre-trained diffusion model. We only train the parameters of the new feature prediction module to preserve object identity and increase generalization. All parameters of the pre-trained model remain frozen, thus keeping our method computationally and storage efficient.
We build our method on Stable Diffusion-XL [64] and show results on various object categories, such as cars, chairs, motorcycles, teddy bears, and toys. We compare our method with image editing, model customization, and NeRF editing methods. Our method maintains high alignment with the target object and poses while adhering to the user-provided text prompt. We show that directly integrating the 3D object information into the text-to-image model pipeline provides performance gains relative to existing 2D methods. Our method can also be combined with other algorithms [52, 3] for applications like generating objects in different target camera poses while preserving the background, panorama synthesis, or composing multiple concepts. Our code, results, and data are available on our webpage.
2 Related Works
Text-based image synthesis.   Large-scale text-to-image models [77, 69, 22, 34, 99] have become ubiquitous with their capabilities of generating photorealistic images from text prompts. This progress has been driven by the availability of large-scale datasets [80] as well as advancements in model architecture and training objectives [19, 79, 35, 63, 36]. Among them, diffusion models [85, 30, 83] have emerged as a powerful family of models that generate images by gradually denoising Gaussian noise. Their learned priors have been found useful in various applications, such as image editing [52, 29, 102] and 3D creation [65, 47].
Image editing.   One crucial application enabled by the above models is image editing based on text instructions [61]. For example, SDEdit [52] exploits the denoising nature of diffusion models, guiding generation in later denoising timesteps using edit instructions while preserving the input image layout. Various works aim to improve upon this by embedding the input image into the model’s latent space [85, 38, 55, 60], while some use cross-attention and self-attention mechanism for realistic and targeted edits [29, 13, 25, 62, 10]. Recently, several methods train conditional diffusion models to follow user edit instructions or spatial controls [102, 9]. However, existing methods primarily focus on changing style and appearance, while our work enables both viewpoint and appearance control.
Model customization.  While pre-trained models are trained to generate common objects, users often wish to synthesize images with concepts from their own lives. This has given rise to the emerging technique of model personalization or customization [74, 23, 45]. These methods aim at embedding a new concept, e.g., pet dog, toy, personal car, person, etc., into the output space of text-to-image models. This enables generating new images of the concept in unseen scenarios using the text prompt, e.g., my car in a field of sunflowers. To achieve this, various works fine-tune a small subset of model parameters [45, 26, 32, 89] and/or optimize text token embeddings [23, 92, 104, 1] on the few images of the new concept with different regularizations [74, 45]. More recently, several encoder-based methods have been proposed that train a model on a vast dataset of concept library [81, 2, 24, 94, 46, 90, 75], enabling faster customization during inference. However, none of the existing works allow controlling the camera pose during inference time. In contrast, given the ease of capturing multi-view images of a new concept, in this work, we ask whether we can augment the capabilities of model customization with additional control of the camera pose.
Figure 2: Overview. We propose a model customization method that utilizes $N$ reference images defining the 3D structure of an object $\mathcal{Y}$ (we illustrate with 2 views for simplicity). We modify the diffusion model U-Net with pose-conditioned transformer blocks. Our Pose-conditioned transformer block features a FeatureNeRF module, which aggregates features from the individual viewpoints to target viewpoint $\phi$, as shown in detail in Figure 3. The rendered feature $W_{y}$ is concatenated with the target noisy feature $W_{\mathbf{x}}$ and projected to the original channel dimension. We use the diffusion U-Net itself to extract features of reference images, as shown in the top row. We only fine-tune the new parameters in linear projection layer $l$ and FeatureNerF in $F_{\text{pose}}$ blocks.
View synthesis.   Novel view synthesis aims to render a scene from unseen camera poses, given multi-view images. Recently, the success of volumetric rendering-based approaches like NeRF [54] have led to numerous follow-up works with better quality [4, 5], faster speed [57, 14], and fewer training views [98, 58, 18, 86]. Recent works learn generative models with large-scale multi-view data to learn generalizable representations for novel view synthesis [49, 78, 106, 11, 95, 50]. While our work draws motivation from this line of research, our goal differs - we aim to enable 3D control in text-to-image personalization, rather than capturing real scenes. Recently, Cheng et al. [17] and Höllein et al. [31] propose adding camera pose in text-to-image diffusion models, while we focus on model customization.
3D editing.   Loosely related to our work, many works have been proposed for inserting and manipulating 3D objects within 2D real photographs, using classic geometry-based approaches [37, 15, 41] or recent generative modeling techniques [96, 103, 53]. Instead of editing a single image, our work aims to “edit” the model weights of a pre-trained 2D diffusion model. Many recent works edit [27, 20] or generate [68, 88, 82] a 3D scene given a text prompt or an image. These methods focus on ensuring the multi-view consistency of the scene. Unlike these, we do not aim to edit a 3D multi-view consistent scene, but instead provide additional camera pose control for the new object when customizing text-to-image models.
3 Method
Given multi-view images of an object, we aim to embed the object in the text-to-image diffusion model. We construct our method in order to allow the generation of new variations of the object through text prompts while providing control of the camera pose of the object. Our approach involves fine-tuning a pre-trained text-to-image diffusion model while conditioning it on a 3D representation of the object learned in the diffusion model’s feature space. In this section, we briefly overview the diffusion model and then explain our method in detail.
3.1 Diffusion Models
Diffusion models [83, 30] are a class of generative models that sample images by iterative denoising of a random Gaussian distribution. The training of the diffusion model consists of a forward Markov process, where real data ${\mathbf{x}}_{0}$ is gradually transformed to random noise ${\mathbf{x}}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ by sequentially adding Gaussian perturbations in $T$ timesteps, i.e., ${\mathbf{x}}_{t}=\sqrt{\alpha_{t}}{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$. The model is trained to learn the backward process, i.e.,
$p_{\theta}({\mathbf{x}}_{0}|\mathbf{c})=\int\Bigr{[}p_{\theta}({\mathbf{x}}_{T})\prod p_{\theta}^{t}({\mathbf{x}}_{t-1}|{\mathbf{x}}_{t},\mathbf{c})\Bigr{]}d{\mathbf{x}}_{1:T},$ (1)
The training objective maximizes the variational lower bound, which can be simplified to a simple reconstruction loss:
$\mathbb{E}_{{\mathbf{x}}_{t},t,\mathbf{c},\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[w_{t}||\epsilon-\epsilon_{\theta}({\mathbf{x}}_{t},t,\mathbf{c})||],$ (2)
where $\mathbf{c}$ can be any modality to condition the generation process. The model is trained to predict the noise added to create the input noisy image ${\mathbf{x}}_{t}$. During inference, we gradually denoise a random Gaussian noise for a fixed set of timesteps. Various sampling strategies [85, 51, 35] have been proposed to reduce the number of sampling steps $T$ compared to the usual $1000$ timesteps in training.
3.2 Customization with Camera Pose Control
Model customization aims to condition the model on a new concept, given $N$ representative images of the concept $\mathcal{Y}=\{{\mathbf{y}}_{i}\}_{i=1}^{N}$, i.e., to model $p({\mathbf{x}}|\mathcal{Y},\mathbf{c})$ with text prompt $\mathbf{c}$. In contrast, we aim to additionally condition the model on camera pose, allowing more control in the generation process. Thus, given a set of multi-view images $\{{\mathbf{y}}_{i}\}_{i=1}^{N}$ and the corresponding camera poses $\{\pi_{i}\}_{i=1}^{N}$, we seek to achieve a customized text-to-image model corresponding to the object, i.e., our goal is to learn the conditional distribution $p({\mathbf{x}}|\{({\mathbf{y}}_{i},\pi_{i})\}_{i=1}^{N},\mathbf{c},\phi)$, where $\mathbf{c}$ is text prompt and $\phi$ is the target camera pose. To achieve this, we fine-tune a pre-trained text-to-image diffusion model, which models $p({\mathbf{x}}|\mathbf{c})$, with the additional conditioning of target pose and reference views.
Model architecture.   In Figure 2, we show our architecture, with an emphasis on the added pose-conditioned transformer block. To begin, we use the Stable Diffusion-XL (SDXL) [64] as the pre-trained text-to-image diffusion model in our work. It is based on the Latent Diffusion Model (LDM) [72], which is trained in an autoencoder [43] latent space. The diffusion model is a U-Net [73] consisting of encoder, middle, and decoder blocks. Each block consists of a ResNet [28], denoted as $h$, followed by several transformer layers [91]. Each transformer layer consists of a self-attention layer (denoted as $s$), followed by a cross-attention (denoted as $g$) with the text condition, and a feed-forward MLP (denoted as $f$). Given feature map ${\mathbf{z}}$, output of an intermediate ResNet layer $h$ in the U-Net, a standard transformer block performs $F_{\text{standard}}({\mathbf{z}},\mathbf{c})=f(g(s({\mathbf{z}}),\mathbf{c}))$. We modify the transformer layer to incorporate pose conditioning.
Pose-conditioned transformer block.   To condition the model on the 3D structure of the object, we modify the transformer block to be a pose-conditioned transformer block $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\},\mathbf{c},\phi)$, where ${\mathbf{z}}_{0}$ is the feature from the main branch, $\{{\mathbf{z}}_{i}\}$ are intermediate feature maps corresponding to multi-view images, and $\{\pi_{i}\}$ and $\phi$ are reference and target camera poses. To condition on the multi-view images, we learn a radiance field conditioned on reference view features in a feed-forward manner [98]. We extract features $\{{\mathbf{W}}_{i}\in\mathbb{R}^{H\times W\times C}\}$ from $\{{\mathbf{z}}_{i}\}$. We use components of pre-trained U-Net itself, $F_{\text{standard}}$, to extract these features and render them into a target pose $\phi$ using a FeatureNeRF function to obtain 2D feature map ${\mathbf{W}}_{y}$.
$\begin{split}{\mathbf{W}}_{i}=F_{\text{standard}}({\mathbf{z}}_{i},\mathbf{c}),\hskip 5.69054pt{\mathbf{W}}_{y}=\text{FeatureNeRF}(\{{\mathbf{W}}_{i},\pi_{i}\},\mathbf{c},\phi)\end{split}$ (3)
FeatureNeRF.   Here, we describe the aggregation of individual 2D features ${\mathbf{W}}_{i}$ with 3D poses $\pi_{i}$ into a feature map ${\mathbf{W}}_{y}$ from pose $\phi$. Rather than learning NeRF in a feature space [97, 40], our focus is on learning 3D features that the 2D diffusion model can use. From target viewpoint $\phi$, for each point $\mathbf{p}$ on a target ray with direction $\mathbf{d}$, we project the point on the image plane of given views $\pi_{i}$ and denote projected locations as $\pi_{i}^{\mathbf{p}}$. We then sample from these coordinates on feature map ${\mathbf{W}}_{i}$, predict a feature for the 3D point, and aggregate them with function $\psi$:
$\mathbf{V}_{i}=$ $\text{MLP}(\text{Sample}({\mathbf{W}}_{i};\pi_{i}^{\mathbf{p}}),\gamma(\mathbf{d}),\gamma(\mathbf{p})),\;{i=1,...,N}$ (4)
$\bar{\mathbf{V}}=\psi(\mathbf{V}_{1},...,\mathbf{V}_{N}),$
Figure 3: FeatureNeRF block. We predict volumetric features $\bar{\mathbf{V}}$ for each 3D point in the grid using reference features $\{W_{i}\}$ (Eqn. 4). Given this feature, we predict the density $\sigma$ and color $rgb$ using a 2-layer MLP and use the predicted density to render $\hat{\mathbf{V}}$, which has been updated with text cross-attention $g$. The predicted $rgb$ is only used to calculate reconstruction loss during training.
where $\gamma$ is the frequency encoding. We use the weighted average [71] as the aggregation function $\psi$, where a linear layer predicts the weights based on $\mathbf{V}_{i}$, $\pi_{i}$, and target pose $\phi$. For each reference view, $\mathbf{d}$ and $\mathbf{p}$ are first transformed in the view coordinate space. We then predict the density and color of the point using a linear layer:
$(\sigma,\mathbf{C})=\text{MLP}(\bar{\mathbf{V}}).$ (5)
We also update the aggregated feature with text condition $\mathbf{c}$ using cross-attention:
$\hat{\mathbf{V}}=\text{CrossAttn}(\bar{\mathbf{V}},\mathbf{c}),$ (6)
and render this updated intermediate feature using the predicted density $\sigma$:
${\mathbf{W}}_{y}(r)=\sum_{j=1}^{N_{f}}T_{j}(1-\exp(-\sigma_{j}\delta_{j}))\hat{\mathbf{V}}_{j},$ (7)
where $r$ is the target ray, $\hat{\mathbf{V}}$ is the aggregated feature of the point in the ray, $\sigma$ is the predicted density of that point, $N_{f}$ is the number of sampled points along the ray between the near and far plane of the camera, and $T_{j}=\exp(-\sum_{k=1}^{j-1}\sigma_{k}\delta_{k})$ handles occlusion until that point.
Conditioning.   To process the main target branch, we extract the intermediate 2D feature map after the self $s$ and cross-attention layers $g$, i.e., ${\mathbf{W}}_{{\mathbf{x}}}=g(s({\mathbf{z}}_{0}),\mathbf{c})$. We concatenate ${\mathbf{W}}_{{\mathbf{x}}}$ with the rendered features ${\mathbf{W}}_{y}$ and then project back into the original feature dimension using a linear layer. Thus, the output of the modified transformer layer is:
$F_{\text{pose}}=f(l({\mathbf{W}}_{y}\oplus{\mathbf{W}}_{{\mathbf{x}}})),$ (8)
where $l$ is a learnable weight matrix, which projects the feature into the space to be processed by the feedforward layer $f$. We initialize $l$ such that the contribution from ${\mathbf{W}}_{y}$ is zero at the start of training.
Training loss.   Our training objective includes learning 3D consistent FeatureNeRF modules, which can contribute to the final goal of reconstructing the target concept in diffusion models output space. Thus, we fine-tune the model using the sum of training losses corresponding to FeatureNeRF and the default diffusion model reconstruction loss:
$\mathcal{L}_{\text{diffusion}}=\sum_{r}Mw_{t}||\epsilon-\epsilon_{\theta}({\mathbf{x}}_{t},t,\mathbf{c})||,$ (9)
where $M$ is the object mask, with the reconstruction loss being calculated only in the object mask region. The losses corresponding to FeatureNeRF consist of RGB reconstruction loss:
$\mathcal{L}_{\text{rgb}}=\sum_{r}||M(r)(\mathbf{C}_{gt}(r)-\sum_{j=1}^{N_{f}}T_{j}(1-\exp(-\sigma_{j}\delta_{j}))\mathbf{C})||,$ (10)
and two mask-based losses – (1) silhouette loss [70] $\mathcal{L}_{\text{s}}$ which forces the rendered opacity to be similar to object mask, and (2) background suppression loss [6, 7] $\mathcal{L}_{\text{bg}}$ which enforces the density of all background rays to be zero, as we only wish to model the object.
$\mathcal{L}_{\text{s}}$ $=\sum_{r}||M(r)-\sum_{j=1}^{N_{f}}T_{j}(1-\exp(-\sigma_{j}\delta_{j}))||$ (11)
$\mathcal{L}_{\text{bg}}=\sum_{r}(1-M(r))\sum_{j=1}^{N_{f}}||(1-\exp(-\sigma_{j}\delta_{j}))||,$
Thus, the final training loss is:
$\mathcal{L}=\mathcal{L}_{\text{diffusion}}+\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{bg}}\mathcal{L}_{\text{bg}}+\lambda_{\text{s}}\mathcal{L}_{\text{s}},$ (12)
where $M$ is the object mask and $\lambda_{i}$ are hyperparameters to control the rendering quality of intermediate images vs the final denoised image. We keep $\lambda_{i}$ fixed across all experiments. We assume access to the object’s mask in the image, which is used to calculate the above losses. The three losses corresponding to FeatureNeRF are averaged across all pose-conditioned transformer layers.
Figure 4: Qualitative comparison. Given a particular target pose, we show the qualitative comparison of our method with (1) Image editing methods SDEdit, InstructPix2Pix, and LEDITS++ which edit a NeRF rendered image from the input pose, (2) Vica-NeRF, a 3D editing method that trains a NeRF model for each input prompt, and (3) LoRA + Camera pose, our proposed baseline where we concatenate camera pose information to text embeddings during LoRA fine-tuning. Our method performs on par or better in keeping the target identity and poses while incorporating the new text prompt—e.g., putting a picnic table next to the SUV car ($1^{\text{st}}$ column)—and following multiple text conditions—e.g., turning the chair red and placing it in a white room ($3^{\text{rd}}$ column). V token is used only in ours and the LoRA + Camera pose method. Ground truth rendering from the given pose is shown as an inset in the first three rows. We show more sample comparisons in Figure 15 of Appendix.
Figure 5: Qualitative samples with varying pose. Our method’s results with different text prompts and target poses as conditions. Our method learns the identity of custom objects while allowing the user to control the camera pose and text prompt for generating the object in new contexts, e.g., changing the background scene or object color and shape. In each row, the images were generated with the same seed while changing the camera pose around the object in a turntable manner. Figure 16 in the Appendix shows more such samples. Note that each image in a row is independently generated. We do not aim to generate multi-view consistent scenes.
Inference.   During inference, to balance the text vs. reference view conditions in the final generated image, we combine text and image guidance [9] as shown below:
$\hat{\epsilon}_{\theta}({\mathbf{x}}_{t},I=\{{\mathbf{y}}_{i},\pi_{i}\}_{i=1}^{N},$ $\mathbf{c})=\epsilon_{\theta}({\mathbf{x}}_{t},\varnothing,\varnothing)$ (13)
$+\lambda_{I}(\epsilon_{\theta}({\mathbf{x}}_{t},I,\varnothing)-\epsilon_{\theta}({\mathbf{x}}_{t},\varnothing,\varnothing))$
$+\lambda_{c}(\epsilon_{\theta}({\mathbf{x}}_{t},I,\mathbf{c})-\epsilon_{\theta}({\mathbf{x}}_{t},I,\varnothing)),$
where $\lambda_{I}$ is the image guidance scale and $\lambda_{c}$ is the text guidance scale. Increasing the image guidance scale increases the generated image’s similarity to the reference images. Increasing the text guidance scale increases the generated image’s consistency with the text prompt.
Training details.   During training, we sample the $N$ views equidistant from each other and use the first as the target viewpoint and the others as references. We modify $12$ transformer layers with pose conditioning out of $70$ transformer layers in Stable Diffusion-XL. For rendering, we sample $24$ points along the ray. The new concept is described as “V category”, with V as a trainable token embedding [45, 23]. Furthermore, to reduce overfitting [74], we use generated images of the same category, such as random car images with ChatGPT-generated captions [12]. These images are randomly sampled $25\%$ of the time during training. We also drop the text prompt with $10\%$ probability to be able to use classifier-free guidance. We provide more implementation details in Appendix C.
4 Experiments
Dataset.   For our experiments, we select concepts from the Common Objects in 3D (CO3Dv2) dataset [71], commonly used for novel view synthesis, and NAVI [33]. Specifically, we select four categories with three instances from the CO3Dv2 dataset- car, chair, teddy bear, and motorcycle—as each instance is uniquely identifiable for these categories. From the NAVI dataset, we select two unique toy-like concepts. We use the camera pose provided in the dataset for the multi-view images. A representative image of each concept is shown in Figure 14, Appendix A. For each instance, we sample $\sim 100$ images and use half for training and half for evaluation. The camera poses are normalized such that the mean of camera location is the origin, and the first camera is at unit norm [100].
Method Text Alignment Image Alignment Photorealism
[…]
Table 1: Human preference evaluation. Our method is preferred over all baselines for text alignment, image alignment to target concept, and photorealism except LoRA + Camera pose, which overfits the training images, as also shown in Figure 4.
Method Angular error Camera center error
Ours 14.19 0.080
LoRA + Camera pose 41.14 0.305
Table 2: Camera pose accuracy in generated images by ours and the LoRA + Camera pose baseline method. We observe that the baseline usually overfits to training images and does not respect the target pose with new text prompts.
Baselines.   While no prior method targets our exact task, we use three types of related baselines – (1) 2D image editing methods, which aim to preserve details of the input image and thus keep the object in the same pose as the input image. This includes three recent and publicly available methods: LEDITS++ [8], InstructPix2Pix [9], and SDEdit [52] with Stable Diffusion -1.5 (and SDXL in Appendix A). As image editing methods by themselves do not support camera viewpoint manipulation, we first render a NeRF model [87] in the target pose and then edit the rendered image. (2) Customization-based method, LoRA+Camera pose, where we modify LoRA [76, 32] by concatenating the camera pose information to the text embeddings, following recent work Zero-1-to-3 [49]. (3) VICA-NeRF [20], a 3D editing method that trains a NeRF for each new text prompt. In Appendix C, we provide more details on implementation and hyperparameters for each baseline.
Evaluation metrics.   To create an evaluation set, we generate $16$ prompts per object category using ChatGPT [12]. We instruct ChatGPT to propose four types of prompts: scene change, color change, object composition, and shape change. We then manually inspect them to remove implausible or overly complicated text prompts [93]. Table 5 in Appendix B lists all the evaluation prompts.
For a quantitative comparison, we primarily use a pairwise human preference study. We compare our method against each baseline to measure image alignment to target concept, alignment to input text prompt, and photorealism of generated images. In total, we collect $\sim 1000$ responses per pairwise study using Amazon Mechanical Turk. We also show the performance of our method and baselines on other standard metrics like CLIP Score [67] and DINOv2 [59] image similarity [74] to measure the text- and image-alignment.
To measure whether the object in generated images corresponds to the input camera pose for our method and the LoRA + Camera pose baseline, we use a pretrained model, RayDiffusion [101], to predict the poses from generated images and calculate its error relative to the ground truth camera poses. More details about evaluation are provided in Appendix B.
Figure 6: Quantitative comparison. We show CLIP scores (higher is better) vs. DINO-v2 scores (higher is better). We plot the performance of each method on each category and the overall mean and standard error (highlighted). Our method results in higher CLIP text alignment while maintaining visual similarity to target concepts, as indicated by DINO-v2 scores. The text alignment of our method compared to SDEdit and InstructPix2Pix is only marginally better as these methods incorporate the text prompt but at the cost of photorealism, as we show in Table 1.
4.1 Results
Figure 7: Extrapolating camera pose from training views. Our method can generalize to different camera poses, including viewpoints not within the training distribution. Top left: We vary the focal length from ×$\times 0.8$ to ×$\times 1.4$ of the original focal length. Top right: We vary the camera position towards the image plane along the $z$ axis. Bottom row: We vary the camera position along the horizontal and vertical axis.
Generation quality and adherence.   First, we measure the quality of the generation – adherence to the text prompt, the identity preservation to the customized objects, and photorealism – irrespective of the camera pose. For the comparison, we generate $18$ images per prompt on $6$ target camera poses, totaling $288$ images per concept. Table 1 shows the pairwise human preference for our method vs. baselines. Our method is preferred over all baselines except LoRA + Camera pose, which we observe to overfit on training images. Figure 6 shows the CLIP vs. DINO scores for all methods and object categories. Ideally, a method should have both a high CLIP score and a DINO score, but often, there is a trade-off between text- and image alignment. Our method has on-par or better text alignment relative to the baselines while having better image alignment. We observe that image-editing baselines often require careful hyperparameter tuning for each image. We select the best-performing hyperparameters and keep them fixed across all experiments. We use the $\sim 50$ validation camera poses not used during training for evaluation and randomly perturb the camera position or focal length. Figure 11 in Appendix B shows sample training and perturbed validation camera poses for the car object.
Camera pose accuracy.   Previously, we have measured our method purely on image customization benchmarks. Next, we evaluate the camera pose accuracy as well. Table 2 shows the camera pose accuracy of the generated images in terms of mean angular error and camera center error. We observe that LoRA + Camera pose baseline overfits on training images and can fail at generating images in the correct pose with new text prompts during inference. We evaluate this on validation camera poses of concepts from the CO3Dv2 dataset with the camera’s principal axis pointing towards the object at the scene’s center. This is because RayDiffusion has been trained on this setup of the CO3Dv2 dataset and fails on unique objects.
Qualitative comparison.   We show the qualitative comparison of our method with the baselines in Figure 4. As we can see, image-editing-based methods often fail at generating photorealistic results. In the case of LoRA + Camera Pose, we observe that it fails to generalize and overfits to the training views ($5^{\text{th}}$ row Figure 4). Finally, the 3D editing-based method Vica-NeRF maintains 3D consistency but generates blurred images for text prompts that change the background scene. Figure 5 shows samples with different text prompts and target camera poses for our method.
Generalization to novel camera poses.   Since our method learns a 3D radiance field, we can also extrapolate to unseen camera poses at inference time as shown in Figure 7. We generate images while varying the camera distance from the object (scale), focal length, or camera position along the horizontal and vertical axis.
Figure 8: Applications. $1^{\text{st}}$ row: Our method can be combined with other image editing methods as well. We use SDEdit with our method to in-paint the rubber duck in different poses while keeping the same background. $2^{\text{nd}}$ row: We can generate interesting panorama shots by controlling the camera pose of the object in each grid independently. $3^{\text{rd}}$ row: We can also compose the radiance field predicted by FeatureNeRF to control the relative pose while generating multiple instances of the object.
Applications.   Our method can be combined with existing image editing methods as well. Figure 8a shows an example where we use SDEdit [52] to in-paint the object in varying poses while keeping the same background.
We can also generate interesting panoramas using MultiDiffusion [3], where the object’s placement in each grid is controlled by our method, as shown in Figure 8b. Moreover, since we learn a 3D consistent FeatureNeRF for the new concept, we can compose multiple instances of the object in feature space [84], with each instance in a different camera pose. Figure 8c shows an example of two teddy bears facing each other sitting on the armchair. Here, we additionally use DenseDiffusion [42] to modulate the attention maps and guide the generation of each object instance to only appear in the corresponding region predicted by FeatureNeRF. At the same time, the attention maps of the empty region predicted by FeatureNeRF are modulated to match the part of the text prompt describing the image’s background.
Method Text Align. Image Align. Camera-pose Accuracy
CLIP- $\uparrow$ fore- $\uparrow$ back- $\downarrow$ Angular error $\downarrow$ Camera center error $\downarrow$
Ours 0.248 0.471 0.348 14.19 0.080
w/o Eqn. 6 0.250 0.460 0.340 16.08 0.096
w/o $\mathcal{L}_{bg}+\mathcal{L}_{s}$ 0.239 0.471 0.371 11.83 0.068
Table 3: Ablation experiments. Not enriching volumetric features with text cross-attention (Eqn. 6) has an adverse effect on image alignment. Not having mask-based losses (Eqn. 11) leads to overfitting on training images and decreases the text alignment. The worst performing metrics are grayed. Our final method achieves a balance between the input conditions of the target concept, text prompt, and camera pose.
4.2 Ablation
In this section, we perform ablation experiments regarding different components of our method and show its contribution. All ablation studies are done on CO3D-v2 instances with validation camera poses.
Background losses.   When removing the silhouette and background loss, as explained in Eqn. 11 from training, we observe a decrease in text alignment and overfitting on training images as shown in Table 3. Figure 12 in Appendix A shows qualitatively that the model generates images with backgrounds more similar to the training views. This is also reflected by the higher similarity between generated images and background regions of the training images ($3^{\text{rd}}$ column Table 3) compared to our final method.
Text cross-attention in FeatureNeRF.   We also enrich the 3D learned features with text cross-attention as shown in Eqn. 6. We perform the ablation experiment of removing this component from the module. Table 3 shows that this leads to a drop in image alignment with the target concept. Thus, cross-attention with text in the volumetric feature space helps the module learn the target concept better.
Figure 9: Limitations. Our method can fail when extrapolating camera poses far from the training image camera poses, e.g., changing focal length (top left) or translating camera s.t. the object is not in the center (top right) as the pre-trained model is often biased towards generating the object in the center. Also, it can fail to follow the input text prompt or the exact camera pose when multiple objects are composed in a scene (bottom row).
5 Discussion and Limitations
We introduce a new task of customizing text-to-image models with camera viewpoint control. Our method jointly learns 3D feature prediction modules and adapts 2D diffusion attention modules to be conditioned on these features. This enables synthesizing the object with new text prompts and precise object pose control. While our method outperforms existing image editing and model customization approaches, it still has several limitations, as we discuss below.
Limitations.   As we show in Figure 9, our method struggles at generalizing to extreme camera poses that were not seen during training and resorts to either changing the object identity or generating the object in a seen pose. We expect this to improve by adding more camera pose variations during training. An additional case when our method struggles to follow the camera pose is when the text prompt adds multiple new objects. We hypothesize that in such challenging scenarios, the model is biased towards generating front views, often seen in training. Also, our proposed pose-conditioning module is trained with a finetuning-based method. Exploring pose-conditioning in a zero-shot customization and editing methods [16, 24] may help reduce the time and computation as well. Here, we focus on enabling camera view control while generating rigid objects. Future work includes extending this conditioning to handle dynamic objects that change the pose in between reference views. One potential way to address this is using a representation based on dynamic and non-rigid NeRF methods [21, 66, 84].
Acknowledgment.   We are thankful to Kangle Deng, Sheng-Yu Wang, and Gaurav Parmar for their helpful comments and discussion and to Sean Liu, Ruihan Gao, Yufei Ye, and Bharath Raj for proofreading the draft. This work was partly done by Nupur Kumari during the Adobe internship. The work is partly supported by Adobe Research, the Packard Fellowship, the Amazon Faculty Research Award, and NSF IIS-2239076. Grace Su is supported by the NSF Graduate Research Fellowship (Grant No. DGE2140739).
Appendix
In Appendix A, we show more results and baseline comparisons. In Appendix B, we provide details regarding the evaluation and human preference study. Finally, in Appendix C, we describe all the implementation details of our method and baselines.
Appendix A Results
More baselines.   We add a comparison to SDEdit with the Stable Diffusion-XL model here, which we found to perform worse or on par with the Stable Diffusion-1.5 version of the model. Figure 10 shows the quantitative comparison.
Per category comparison.   We show text-alignment and image-alignment scores of our method and baselines for each category in Figure 10 while varying the text guidance scale for each method from $5.0$ to $10.0$, except for LEDITS++ [8] which recommends varying the concept editing guidance scale from $10.0$ to $15.0$. For each guidance scale, we generate $288$ images. Our method usually results in a higher CLIP score than the baselines, indicating better text alignment. For all categories, we show the linear fit curve over the different guidance scales, and our method lies at the Pareto-frontal compared to the baseline methods, either performing similarly or better.
Using predicted masks for background losses.   In our main paper, we calculated the background losses (Eqn. 11) using ground truth masks from the dataset during training. We show in Table 4 that replacing these with predicted masks results in similar performance. We use Detic [105] with Segment Anything [44] to predict the object mask given the object category, such as car and teddy bear.
Varying number of views.   In all our experiments, we use $\sim 50$ multi-view images and their ground truth poses for training. Here, we vary the number of multi-view images to $35$ and $20$, respectively, and show its results in Table 4. As the number of views decreases, text alignment remains similar, but the accuracy of camera pose and image alignment gradually decreases. Example generations and their comparison to our main method are shown in Figure 13.
Qualitative comparison.   We show more qualitative comparisons between our method and baselines in Figure 15. Figure 16 shows more samples from varying camera poses for our method.
Comparison to LoRA.   We compare our method to the customization method LoRA [32, 76] on the text and image alignment metrics here. The CLIP and DINO scores for LoRA are $0.275$ and $0.468$, respectively. In comparison, our CLIP and DINO scores are $0.253$ and $0.481$, respectively. Though the CLIP score is marginally lower, our method performs better in preserving the target concept while allowing additional camera pose control for the custom object.
Method Text Alignment Image Alignment Camera-pose Accuracy
[…]
Ours 0.248 0.471 14.19 0.080
w/ predicted mask 0.246 0.475 13.80 0.086
w/ 35 views 0.250 0.439 18.09 0.108
w/ 20 views 0.254 0.448 18.96 0.108
Table 4: Results with predicted masks and variable views. Our method works similarly even when using predicted masks instead of ground truth masks for mask-based losses. When varying the number of views, the performance gradually drops w.r.t. camera pose accuracy and image alignment.
Appendix B Evaluation.
Evaluation text prompts.   As mentioned in the main paper, we used ChatGPT to generate text prompts for evaluating our method and baselines. Table 5 lists all of the prompts. An example instruction given to ChatGPT to generate prompts for the motorcycle category is: “Provide 16 diverse captions for plausible naturally occurring images of a motorcycle. Only follow one of the options given below while generating the captions. Thus, four captions per option. Each caption should have a simple sentence structure. (1) change the background scene of the motorcycle, (2) insert a new object in the scene with the motorcycle (3) change the type of the motorcycle (4) change the color of the motorcycle”.
Evaluation camera pose.   Figure 11 shows sample training and validation camera poses for the car object. To measure the camera pose accuracy, we randomly select $6$ validation camera poses (Figure 11, $2^{\text{nd}}$ column) and generate images using one of the $16$ text prompts. We then use RayDiffusion [101] to predict the camera pose from the $6$ generated images and calculate its error with the target camera pose. The validation camera poses are such that the principal axis of the camera points towards the object. For text- and image-alignment metrics, we use the perturbed validation camera poses (Figure 11, $3^{\text{rd}}$ column).
Human preference study.   We perform a pairwise comparison of our method with each baseline. To measure text alignment, we show the input text prompt and the two images generated by ours and the baseline method, respectively, and ask: “Which image is more consistent with the following text?” For the image alignment, we show $3$-$4$ target images at the top and ask: “Which of the below images is more consistent with the shown target object?” To measure photorealism, we only show the two generated images with the question: “Which of the below images is more photorealistic?”
DINO image alignment metrics.   We use DINOv2 [59] as the pretrained model to measure image alignment. For each generated image, we measure its mean similarity to all the training images of the target concept. We crop the object region in the training images using masks to measure the similarity only with the target concept.
Appendix C Implementation Details
C.1 Our Method
We fine-tune a pretrained Stable Diffusion-XL $32$-bit floating point model with a batch size of $4$. We update the new parameters with a learning rate of $1\times 10^{-4}$. During training, we bias the sampling of time towards later time steps [56] since pose information is more crucial in the early stages of denoising. Training is done for $1600$ iterations, which takes $\sim 45$ minutes on $4$ A100-GPUs with $40GB$ VRAM. At each training step, we sample $5$ views (maximum possible in GPU memory) equidistant from each other and use the first as the target viewpoint and the others as references. We crop the reference images tightly around the object bounding box and modify the camera intrinsics accordingly. We modify $12$ transformer layers with pose-conditioning out of a total of $70$ transformer layers in Stable Diffusion-XL, with $4$ in the encoder, $3$ in the middle, and $5$ in the decoder blocks of the U-Net. Further, in a particular encoder or decoder block, we use the density predicted by previous FeatureNeRF blocks to importance sample the points along the ray for the next FeatureNeRF block $90\%$ of the times. This improves the performance on concepts with thin structuress like chairs. For rendering, we sample $24$ points along the ray. The training hyperparameters, $\lambda_{\text{rgb}}$, $\lambda_{\text{s}}$, $\lambda_{\text{bg}}$ are set to the to $5$, $10$, and $10$ for all experiments.
During inference, we use an image guidance scale of $3.5$ and a text guidance scale of $7.5$ in Eqn. 13. All images for evaluation are generated with $50$ sampling steps using the default Euler scheduler [35]. With these settings, the wallclock runtime to generate one image given $8$ reference views is $10$ seconds. We cache the reference view features for each customized model. For comparison, it takes $6$ seconds to generate an image given no reference views.
Figure 10: CLIP vs. DINO scores, separated by object category. For each category, our method achieves higher or the same text alignment compared to baselines while having on-par image alignment.)
Figure 11: Sample training and validation (+perturbed) views. We show sample training and validation views that we use in our method. For quantitative evaluation, we perturb the location and focal length of the validation camera poses to create the final set of evaluation target camera poses.
Figure 12: Role of mask-based loss in FeatureNeRF. Not having silhouette and background losses results in the model overfitting to the background features of the training image, e.g., the trees in the background.
Figure 13: Number of training views. As we decrease the number of training views, some target camera poses are not respected in the output generations, specifically back-facing viewpoints. We observe that the model has a bias towards generating front-facing objects.
C.2 Baselines
For 2D image editing methods, we first render the image given the target camera pose using a trained NeRF model [87]. To remove floater artifacts outside of the object mask in the rendered image, we run a pre-processing step using the SDEdit Stable Diffusion-XL denoising ensemble [64] with negative prompts “blurry, blur”. The inpainted image is then passed to the 2D image editing method as the input. For each baseline, we calculate metrics at $5$ different guidance scales and keep the remaining hyperparameters the same across all object categories as much as possible.
SDEdit [52].   This uses the forward process to create a noisy image at some intermediate timestep and runs the backward denoising process with the new text prompt. For the optimal strength value, we ran a grid search from $0.5$ to $0.8$. We report our evaluation metrics on the best-performing strength of $0.5$ (highest average text alignment to image alignment ratio). We run all inference in float16161616. For SDEdit with Stable Diffusion-1.5, we use the recommended PNDM sampler [48] and set the number of inference steps to $50$. With a strength of $0.5$, the number of denoising steps run is $25$ and takes $1$ second. In the case of Stable Diffusion-XL, we apply the base model and refiner model as an ensemble of expert denoisers [64]. We run the base model with the default Euler scheduler [35] for $15$ steps at strength $0.5$ and the refiner for $5$ more steps. With these settings, the wallclock runtime to generate one image is $5$ seconds.
LEDITS++ [8].   This is a more recent method that proposes a new inversion technique to embed the image in latent space. It then constructs semantically grounded masks from the U-Net’s cross-attention layers and noise estimates to constrain the edit regions corresponding to the new text prompt. The user must adjust the LEDITS++ hyperparameter values according to whether they wish to modify large or small regions of the image, e.g., background changes vs. object appearance edits. To find the best hyperparameters, we refer to the recommended values provided in the official implementation. Thus, for prompts that edit the object’s appearance, we keep the target prompt guidance scale at $8$, the edited concept’s threshold at $0.9$, the default number of inference steps at $50$, and the proportion of initial denoising steps skipped to $0.25$. For prompts that change the background significantly, we additionally change the target prompt from the original image’s BLIP caption [46] to the background-changing prompt. For example, the edit in Figure 13 is achieved in LEDITS++ by changing the target prompt from “A teddy bear sitting on a table in a living room” to “A teddybear on the sand at the beach”. On the other hand, to edit the color of the teddy bear to gray using LEDITS++, the target prompt would still be “A teddy bear sitting on a table in a living room” but the concept “A gray teddybear” would be added to guided the generated image in that direction.
InstructPix2Pix [9].   This image-editing technique trains a model to follow editing instructions and can edit a new input image in a feedforward manner. We use the official model released with the paper based on Stable Diffusion-1.5, set the image guidance scale to the suggested value of $1.5$, and use the default Euler scheduler with $50$ inference steps.
ViCA-NeRF [20].   This is a 3D editing method that provides improved multiview consistency and training speed compared to Instruct-NeRF2NeRF [27]. First, the editing stage edits key views from the set of multiview images using InstructPix2Pix, reprojects the editing results to other views using the camera pose and depth information, and blends the edits in the latent diffusion model’s feature space. Second, the NeRF training stage uses the edited multiview images to optimize the unedited NeRF and produce a 3D-edited NeRF. We use the official implementation’s hyperparameters of text guidance scale $7.5$ and image guidance scale $1.5$. For concepts from the NAVI dataset, we select images from a single scene. For our method, we use images from multiple scenes since we only model the foreground object.
LoRA [32].   This is a 2D customization method that trains low-rank adapters in linear layers of the diffusion model U-Net when fine-tuning on the new custom concept. We fine-tune Stable Diffusion-XL using the LoRA fine-tuning technique with rank $64$ adapters added to all the linear layers in the attention blocks. We use the recommended learning rate of $1\times 10^{-4}$ with batch-size $16$ and train for $1000$ iterations. We sample the regularization images as in our method $25\%$ of the times. For training, we use the text prompt photo of a V {category} , where V is a fixed token.
LoRA + Camera pose.   We modify the LoRA customization method to include the camera pose condition and text prompt. To achieve this, in every cross-attention layer, we concatenate the flattened camera projection matrix with the text transformer output along the feature dimension and use a one-layer MLP to project it back to the original dimension. The one-layer MLP is trained along with LoRA adapter modules, with all other hyperparameters kept the same as the above LoRA baseline. To match our method, we also biased the sampling of time towards later time steps. Like ours, the reconstruction loss is only applied in the masked region. We train the model for a total of $2000$ iterations.
Object Category Evaluation Prompt
Car A car parked by a snowy mountain range.
A car on a bridge over a calm river.
A car in front of an old, brick train station.
A car beside a field of blooming sunflowers.
A car with a bike rack on top.
A car next to a picnic table in a park.
A car with a guitar leaning against it.
A car with a kayak mounted on the roof.
A minivan car outside a school, during pickup time.
A convertible car near a coastal boardwalk.
A jeep car on a rugged dirt road.
A volkswagon beetle car in front of a luxury hotel.
A red car in a mall parking lot.
A yellow car at a gas station.
A green car in a driveway, next to a house.
A black car in a busy city street.
Chair A chair on a balcony overlooking the city skyline.
A chair in a garden surrounded by flowers.
A chair on a beach.
A chair in a library next to a bookshelf.
A chair with a plush toy sitting on it.
A chair beside a guitar on a stand.
A chair next to a potted plant.
A chair with a colorful cushion on it.
A rocking chair on a porch.
An office chair in a home study.
A folding chair at a camping site.
A high chair in a kitchen.
A red chair in a white room.
A black chair in a classroom.
A green chair in a café.
A yellow chair in a playroom.
Teddybear A teddybear on a park bench under trees.
A teddybear at a window with raindrops outside.
A teddybear on the sand at the beach.
A teddybear on a cozy armchair by a fireplace.
A teddybear with a stack of children’s books on the side.
A teddybear next to a birthday cake with candles.
A teddybear with a small toy car.
A teddybear holding a heart-shaped balloon.
A teddybear in a pink barbie costume.
A large teddybear in a batman costume.
A teddybear dressed as a construction worker.
A teddybear in a superhero costume.
A pink teddybear on a shelf.
A brown teddybear on a blanket.
A white teddybear.
A gray teddybear.
Motorcycle A motorcycle parked on a city street at night.
A motorcycle beside a calm lake.
A motorcycle on a mountain road with a scenic view.
A motorcycle in front of a graffiti-covered urban wall.
A motorcycle with a guitar strapped to the back.
A motorcycle next to a camping tent.
A golden retriever riding motorcycle.
A cat riding motorcycle.
A cruiser motorcycle in a parking lot.
A scooter like motorcycle.
A vintage style motorcycle.
A dirt bike motorcycle on a trail in the woods.
A red motorcycle in a garage.
A green motorcycle.
A blue motorcycle.
A silver motorcycle.
Toy Toy on a sandy beach, with waves crashing in the background
A toy sitting in a grassy field, surrounded by wildflowers.
A toy on a rocky mountain top, overlooking the valley below.
A toy in a dense jungle.
A toy with a tiny book placed beside it on a wooden table.
A toy floating next to a colorful beach ball in a bathtub.
A toy with a small globe resting next to it.
A toy and an umbrella in a cozy living room.
A plush toy on a sunny windowsill.
A wooden toy.
An origami of toy.
A clay figurine of toy.
A bright red toy.
A deep blue toy on a bed.
A vivid green toy.
A neon pink toy.
Table 5: Evaluation prompts. Here, we list the final prompts that were used for evaluation for the car, chair, teddy bear, motorcycle, and toy categories.
Figure 14: Target concepts. Sample images from the $14$ target concepts used as the dataset for evaluating our method.
Figure 15: Additional qualitative comparison of our method with baselines, given various prompts and target pose conditions.
Figure 16: Additional qualitative samples of our method while varying the camera pose condition for the custom object.

### v2

@@
-Customizing Text-to-Image Diffusion with Camera Viewpoint Control
+Customizing Text-to-Image Diffusion with Object Viewpoint Control
 Abstract
 * indicates equal contribution
-Model customization introduces new concepts to existing text-to-image models, enabling the generation of the new concept in novel contexts.
-However, such methods lack accurate camera view control w.r.t the object, and users must resort to prompt engineering (e.g., adding “top-view”) to achieve coarse view control.
-In this work, we introduce a new task – enabling explicit control of camera viewpoint for model customization.
-This allows us to modify object properties amongst various background scenes via text prompts, all while incorporating the target camera pose as additional control.
-This new task presents significant challenges in merging a 3D representation from the multi-view images of the new concept with a general, 2D text-to-image model.
-To bridge this gap, we propose to condition the 2D diffusion process on rendered, view-dependent features of the new object.
-During training, we jointly adapt the 2D diffusion modules and 3D feature predictions to reconstruct the object’s appearance and geometry while reducing overfitting to the input multi-view images.
-Our method outperforms existing image editing and model personalization baselines in preserving the custom object’s identity while following the input text prompt and the object’s camera pose.
-Figure 1: Given multi-view images of a new object (left), we create a customized text-to-image diffusion model with camera pose control.
-The customized model can synthesize the object (denoted as V ¡category name¿) in novel appearances and scenes, such as A green V car, or A beetle-like V car, while allowing users to specify the target camera pose for the object.
-We can generate panorama images while controlling the camera pose for each object or compose multiple concepts using MultiDiffusion [3] with our model.
+Model customization introduces new concepts to existing text-to-image models, enabling the generation of these new concepts/objects in novel contexts.
+However, such methods lack accurate camera view control with respect to the new object, and users must resort to prompt engineering (e.g., adding “top-view”) to achieve coarse view control.
+In this work, we introduce a new task – enabling explicit control of the object viewpoint in the customization of text-to-image diffusion models.
+This allows us to modify the custom object’s properties and generate it in various background scenes via text prompts, all while incorporating the object viewpoint as an additional control.
+This new task presents significant challenges, as one must harmoniously merge a 3D representation from the multi-view images with the 2D pre-trained model.
+To bridge this gap, we propose to condition the diffusion process on the 3D object features rendered from the target viewpoint.
+During training, we fine-tune the 3D feature prediction modules to reconstruct the object’s appearance and geometry, while reducing overfitting to the input multi-view images.
+Our method outperforms existing image editing and model customization baselines in preserving the custom object’s identity while following the target object viewpoint and the text prompt.
+Figure 1: Given multi-view images of a new object (left), denoted as V ¡category name¿, we create a customized text-to-image diffusion model with object viewpoint control.
+The customized model allows users to specify the target viewpoint for the object while synthesizing it in novel appearances and scenes, such as A green V car, or A beetle-like V car.
+We can also generate panorama images or compose multiple concepts while controlling each object’s viewpoint by using MultiDiffusion [4] with our model.
 1 Introduction
 Recently, we have witnessed an explosion of works on customizing text-to-image models [74, 25, 47, 18].
-They allow us to quickly acquire visual concepts, such as personal objects and favorite places, and reimagine them with new environments and attributes.
-For instance, we can customize a model on our Teddy bear and prompt it with “Teddy bear on a bench in the park.” Unfortunately, customization methods lack precise camera pose control, as existing diffusion models are trained purely on 2D images without ground truth camera poses.
+Such methods enable a model to quickly acquire visual concepts, such as personal objects and favorite places, and re-imagine them with new environments and attributes.
+For instance, we can customize a model on our teddy bear and prompt it with “Teddy bear on a bench in the park.” However, these methods lack precise viewpoint control, as the pre-trained model is trained purely on 2D images without ground truth camera poses.
 As a result, users often rely on text prompts such as “front-facing” or “side-facing”, a tedious and unwieldy process to control views.
-What if we wish to synthesize a new object, e.g., the Teddy bear in Figure 1, in a different context while controlling its pose?
-In this work, we introduce a new task: given multi-view images of the object, we customize a text-to-image model while enabling precise control of the new object’s camera pose.
-During inference, our method offers the flexibility of conditioning the generation process on both a target pose and a text prompt.
+What if a user wishes to control the custom object’s viewpoint while synthesizing it in a different context, e.g., the car in Figure 1?
+In this work, we introduce a new task: given multi-view images of an object, we customize a text-to-image model while enabling control of the object’s viewpoint.
+During inference, our method offers the flexibility of conditioning the generation process on both a target viewpoint and a text prompt.
 Neural rendering methods have allowed us to accurately control the 3D viewpoint of an existing scene, given multi-view images [41, 59, 6, 5].
-Similarly, we seek to imagine the object from novel viewpoints but in a new scene.
+Similarly, we seek to imagine the object from novel views but in a new context.
 However, as pre-trained diffusion models, such as Latent Diffusion models [72], are built upon a purely 2D representation, connecting the 3D neural representation of the object to the 2D internal features of the diffusion model remains challenging.
-In this work, we propose a new method, CustomDiffusion360, to bridge the gap between 3D neural capture and 2D text-to-image diffusion models by providing additional camera pose control w.r.t.
-the new custom object in 2D text-to-image models.
-More concretely, given multi-view images of an object, we learn to predict neural feature fields in the intermediate feature spaces of the diffusion model U-Net.
-To condition the generation process on a target pose, we render the features of the predicted feature fields and then fuse them with the target pose’s noisy features.
-This new conditioning module is added to a subset of transformer layers in the pre-trained diffusion model.
-We only train the parameters of the new feature prediction module to preserve object identity and increase generalization.
-All parameters of the pre-trained model remain frozen, thus keeping our method computationally and storage efficient.
+In this work, we introduce CustomDiffusion360, a new method to bridge the gap between 3D neural capture and 2D text-to-image diffusion models by providing viewpoint control for custom objects.
+More concretely, given multi-view images of an object, we introduce FeatureNeRF blocks in the diffusion model U-Net’s intermediate feature spaces to learn view-dependent features.
+To condition the generation process on a target viewpoint, we render the FeatureNeRF output from this viewpoint and merge it with the diffusion features using linear projection layers.
+We only train the new linear projection layers and FeatureNeRF blocks, added to a subset of transformer layers, to preserve object identity while maintaining generalization.
+The pre-trained model’s parameters remain frozen, thus keeping our method computationally and storage efficient.
 We build our method on Stable Diffusion-XL [65] and show results on various object categories, such as cars, chairs, motorcycles, teddy bears, and toys.
-We compare our method with image editing, model customization, and NeRF editing methods.
-Our method maintains high alignment with the target object and poses while adhering to the user-provided text prompt.
-We show that directly integrating the 3D object information into the text-to-image model pipeline provides performance gains relative to existing 2D methods.
-Our method can also be combined with other algorithms [52, 3] for applications like generating objects in different target camera poses while preserving the background, panorama synthesis, or composing multiple concepts.
-Our code, results, and data are available on our webpage.
+We compare our approach with image editing [53, 10], model customization [34], and NeRF editing methods [22].
+Our method achieves high alignment with the custom object’s identity and target viewpoint while adhering to the user-provided text prompt.
+We show that integrating the 3D object information into the text-to-image model, as done by our method, enhances performance over 2D and 3D editing baseline methods.
+Additionally, our method can be combined with other algorithms [53, 4] for applications such as object viewpoint adjustment in the same background, panorama synthesis, and object composition.
 2 Related Works
 Text-based image synthesis.
-Large-scale text-to-image models [77, 69, 22, 34, 99] have become ubiquitous with their capabilities of generating photorealistic images from text prompts.
+Large-scale text-to-image models [77, 69, 24, 36, 103] have become ubiquitous for generating photorealistic images from text prompts.
 This progress has been driven by the availability of large-scale datasets [80] as well as advancements in model architecture and training objectives [21, 79, 37, 64, 38].
-Among them, diffusion models [85, 30, 83] have emerged as a powerful family of models that generate images by gradually denoising Gaussian noise.
-Their learned priors have been found useful in various applications, such as image editing [52, 29, 102] and 3D creation [65, 47].
-Image editing.
-One crucial application enabled by the above models is image editing based on text instructions [61].
-For example, SDEdit [52] exploits the denoising nature of diffusion models, guiding generation in later denoising timesteps using edit instructions while preserving the input image layout.
-Various works aim to improve upon this by embedding the input image into the model’s latent space [85, 38, 55, 60], while some use cross-attention and self-attention mechanism for realistic and targeted edits [29, 13, 25, 62, 10].
+Among them, diffusion models [85, 32] have emerged as a powerful family of models that generate images by gradually denoising Gaussian noise.
+Image editing with text-to-image diffusion.
+One of the first works, SDEdit [53], exploited the denoising nature of diffusion models, guiding generation in later denoising timesteps using edit instructions while preserving the input image layout.
+Since then, various works improved upon this by embedding the input image into the model’s latent space [85, 40, 57, 62] or using cross-attention and self-attention mechanisms for realistic and targeted edits [31, 15, 27, 63, 12].
 Recently, several methods train conditional diffusion models to follow user edit instructions or spatial controls [107, 10].
-However, existing methods primarily focus on changing style and appearance, while our work enables both viewpoint and appearance control.
+However, these methods primarily focus on appearance editing, while our work enables both viewpoint and appearance control.
 Model customization.
-While pre-trained models are trained to generate common objects, users often wish to synthesize images with concepts from their own lives.
+While pre-trained models can generate common objects, users often wish to synthesize images with concepts from their own lives.
 This has given rise to the emerging technique of model personalization or customization [74, 25, 47].
-These methods aim at embedding a new concept, e.g., pet dog, toy, personal car, person, etc., into the output space of text-to-image models.
+These methods aim at embedding a new concept, e.g., pet dog, personal car, person, etc., into the output space of text-to-image models.
 This enables generating new images of the concept in unseen scenarios using the text prompt, e.g., my car in a field of sunflowers.
 To achieve this, various works fine-tune a small subset of model parameters [47, 28, 34, 89] and/or optimize text token embeddings [25, 92, 109, 2] on the few images of the new concept with different regularizations [74, 47].
-More recently, several encoder-based methods have been proposed that train a model on a vast dataset of concept library [81, 2, 24, 94, 46, 90, 75], enabling faster customization during inference.
-However, none of the existing works allow controlling the camera pose during inference time.
-In contrast, given the ease of capturing multi-view images of a new concept, in this work, we ask whether we can augment the capabilities of model customization with additional control of the camera pose.
+More recently, encoder-based methods have been proposed that train a model on a vast dataset of concepts [81, 3, 26, 94, 48, 90, 75, 99], enabling faster customization.
+However, to our knowledge, no existing work allows for controlling the viewpoint in model customization.
+Given the ease of capturing multi-view images of a new concept, this work explores augmenting model customization with additional object viewpoint control.
 Figure 2: Overview.
-We propose a model customization method that utilizes $N$ reference images defining the 3D structure of an object $\mathcal{Y}$ (we illustrate with 2 views for simplicity).
+We propose a model customization method that utilizes $N$ reference images defining the 3D structure of an object $\mathcal{Y}$ (we illustrate with $N=2$ views for simplicity).
 We modify the diffusion model U-Net with pose-conditioned transformer blocks.
@@
 Recently, the success of volumetric rendering-based approaches like NeRF [56] have led to numerous follow-up works with better quality [5, 6], faster speed [59, 16], and fewer training views [102, 60, 20, 86].
-Recent works learn generative models with large-scale multi-view data to learn generalizable representations for novel view synthesis [49, 78, 106, 11, 95, 50].
-While our work draws motivation from this line of research, our goal differs - we aim to enable 3D control in text-to-image personalization, rather than capturing real scenes.
+Recent works learn generative models with large-scale multi-view data to learn generalizable representations for novel view synthesis [50, 78, 111, 13, 95, 51, 11].
+While our work draws motivation from this line of research, our goal differs - we aim to enable object viewpoint control in text-to-image personalization, rather than capturing novel views of real scenes.
+Concurrent to our work, ReconFusion [95] also trains a PixelNeRF [102] in the latent space of latent diffusion models for 3D reconstruction.
+Different from this, we learn volumetric features in the intermediate attention layers.
+We also focus on model customization rather than scene reconstruction.
 Recently, Cheng et al.
 [19] and Höllein et al.
-[31] propose adding camera pose in text-to-image diffusion models, while we focus on model customization.
+[33] propose adding camera pose conditioning in text-to-image diffusion models while we focus on model customization.
+CustomNet [104], a concurrent work, also proposes to generate custom objects in a target viewpoint in a zero-shot manner.
+However, it focuses primarily on generating the new object in different backgrounds, whereas our method allows any new text prompt and viewpoint combination as a condition during inference.
 3D editing.
-Loosely related to our work, many works have been proposed for inserting and manipulating 3D objects within 2D real photographs, using classic geometry-based approaches [37, 15, 41] or recent generative modeling techniques [96, 103, 53].
-Instead of editing a single image, our work aims to “edit” the model weights of a pre-trained 2D diffusion model.
-Many recent works edit [27, 20] or generate [68, 88, 82] a 3D scene given a text prompt or an image.
-These methods focus on ensuring the multi-view consistency of the scene.
-Unlike these, we do not aim to edit a 3D multi-view consistent scene, but instead provide additional camera pose control for the new object when customizing text-to-image models.
+Loosely related to our work, many works have been proposed for inserting and manipulating 3D objects within 2D real photographs by editing the image, using classic geometry-based approaches [39, 17, 43] or generative modeling techniques [98, 108, 55, 96, 101].
+Instead of editing a single image, our work aims to “edit” the model weights of a pre-trained diffusion model.
+Another relevant line of work edits [29, 22] or generates [68, 88, 82, 97, 54] a 3D scene given a text prompt or image.
+Unlike these methods, we do not aim to edit/generate a multi-view consistent scene.
+Our goal is to provide additional viewpoint control when customizing text-to-image models.
+This enables specifying the object viewpoint while generating new backgrounds or composing multiple objects.
+Additionally, we show that our method achieves greater photorealism compared to a 3D editing method for this task.
 3 Method
-Given multi-view images of an object, we aim to embed the object in the text-to-image diffusion model.
-We construct our method in order to allow the generation of new variations of the object through text prompts while providing control of the camera pose of the object.
-Our approach involves fine-tuning a pre-trained text-to-image diffusion model while conditioning it on a 3D representation of the object learned in the diffusion model’s feature space.
+Given multi-view images of a custom object, we aim to embed it in the text-to-image diffusion model.
+We construct our method in order to allow the generation of new variations of the object through text prompts while providing control of the object viewpoint.
+Our approach involves fine-tuning the pre-trained model while conditioning it on a 3D representation of the object learned in the diffusion model’s feature space.
 In this section, we briefly overview the diffusion model and then explain our method in detail.
@@
 The model is trained to predict the noise added to create the input noisy image ${\mathbf{x}}_{t}$.
-During inference, we gradually denoise a random Gaussian noise for a fixed set of timesteps.
-Various sampling strategies [85, 51, 35] have been proposed to reduce the number of sampling steps $T$ compared to the usual $1000$ timesteps in training.
-3.2 Customization with Camera Pose Control
-Model customization aims to condition the model on a new concept, given $N$ representative images of the concept $\mathcal{Y}=\{{\mathbf{y}}_{i}\}_{i=1}^{N}$, i.e., to model $p({\mathbf{x}}|\mathcal{Y},\mathbf{c})$ with text prompt $\mathbf{c}$.
-In contrast, we aim to additionally condition the model on camera pose, allowing more control in the generation process.
-Thus, given a set of multi-view images $\{{\mathbf{y}}_{i}\}_{i=1}^{N}$ and the corresponding camera poses $\{\pi_{i}\}_{i=1}^{N}$, we seek to achieve a customized text-to-image model corresponding to the object, i.e., our goal is to learn the conditional distribution $p({\mathbf{x}}|\{({\mathbf{y}}_{i},\pi_{i})\}_{i=1}^{N},\mathbf{c},\phi)$, where $\mathbf{c}$ is text prompt and $\phi$ is the target camera pose.
-To achieve this, we fine-tune a pre-trained text-to-image diffusion model, which models $p({\mathbf{x}}|\mathbf{c})$, with the additional conditioning of target pose and reference views.
+During inference, we gradually denoise a random Gaussian noise over a fixed number of timesteps.
+Various proposed sampling strategies [85, 52, 37] reduce the number of sampling steps compared to the typical $1000$ timesteps in training.
+In our work, we use the Stable Diffusion-XL (SDXL) [65] as the pre-trained text-to-image diffusion model.
+It is based on the Latent Diffusion Model (LDM) [72], which is trained in an autoencoder [45] latent space.
+3.2 Customization with Object Viewpoint Control
+Model customization aims to condition the model on a new object, given $N$ images of the object $\mathcal{Y}=\{{\mathbf{y}}_{i}\}_{i=1}^{N}$, i.e., to model $p({\mathbf{x}}|\mathcal{Y},\mathbf{c})$ with text prompt $\mathbf{c}$.
+In contrast, we additionally condition the model on the object viewpoint, allowing more control in the generation process.
+Thus, given a set of multi-view images $\{{\mathbf{y}}_{i}\}_{i=1}^{N}$ and the corresponding camera poses $\{\pi_{i}\}_{i=1}^{N}$, our goal is to learn the conditional distribution $p({\mathbf{x}}|\{({\mathbf{y}}_{i},\pi_{i})\}_{i=1}^{N},\mathbf{c},\phi)$, where $\mathbf{c}$ is text prompt and $\phi$ is the camera pose corresponding to the target viewpoint.
+To achieve this, we fine-tune a pre-trained text-to-image diffusion model, which models $p({\mathbf{x}}|\mathbf{c})$, with the additional conditioning of camera pose $\phi$ given posed reference images $\{{\mathbf{y}}_{i},\pi_{i}\}_{i=1}^{N}$.
 Model architecture.
-In Figure 2, we show our architecture, with an emphasis on the added pose-conditioned transformer block.
-To begin, we use the Stable Diffusion-XL (SDXL) [64] as the pre-trained text-to-image diffusion model in our work.
-It is based on the Latent Diffusion Model (LDM) [72], which is trained in an autoencoder [43] latent space.
-The diffusion model is a U-Net [73] consisting of encoder, middle, and decoder blocks.
-Each block consists of a ResNet [28], denoted as $h$, followed by several transformer layers [91].
-Each transformer layer consists of a self-attention layer (denoted as $s$), followed by a cross-attention (denoted as $g$) with the text condition, and a feed-forward MLP (denoted as $f$).
-Given feature map ${\mathbf{z}}$, output of an intermediate ResNet layer $h$ in the U-Net, a standard transformer block performs $F_{\text{standard}}({\mathbf{z}},\mathbf{c})=f(g(s({\mathbf{z}}),\mathbf{c}))$.
-We modify the transformer layer to incorporate pose conditioning.
-Pose-conditioned transformer block.
-To condition the model on the 3D structure of the object, we modify the transformer block to be a pose-conditioned transformer block $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\},\mathbf{c},\phi)$, where ${\mathbf{z}}_{0}$ is the feature from the main branch, $\{{\mathbf{z}}_{i}\}$ are intermediate feature maps corresponding to multi-view images, and $\{\pi_{i}\}$ and $\phi$ are reference and target camera poses.
-To condition on the multi-view images, we learn a radiance field conditioned on reference view features in a feed-forward manner [98].
-We extract features $\{{\mathbf{W}}_{i}\in\mathbb{R}^{H\times W\times C}\}$ from $\{{\mathbf{z}}_{i}\}$.
-We use components of pre-trained U-Net itself, $F_{\text{standard}}$, to extract these features and render them into a target pose $\phi$ using a FeatureNeRF function to obtain 2D feature map ${\mathbf{W}}_{y}$.
-$\begin{split}{\mathbf{W}}_{i}=F_{\text{standard}}({\mathbf{z}}_{i},\mathbf{c}),\hskip 5.69054pt{\mathbf{W}}_{y}=\text{FeatureNeRF}(\{{\mathbf{W}}_{i},\pi_{i}\},\mathbf{c},\phi)\end{split}$ (3)
+In Figure 2, we show the overall architecture, with an emphasis on our added pose-conditioning.
+Each block in the diffusion model U-Net [73] consists of a ResNet [30], denoted as $h$, followed by several transformer layers [91].
+Given the output of an intermediate ResNet layer ${\mathbf{z}}$, a standard transformer layer, $F_{\text{standard}}({\mathbf{z}},\mathbf{c})$, consists of a self-attention layer, denoted as $s$, followed by cross-attention with the text prompt, denoted as $g$, and a feed-forward MLP, denoted as $f$.
+We modify a subset of these transformer layers to incorporate pose conditioning as we explain next.
+Pose-conditioned transformer layer.   We denote the pose-
+conditioned transformer layer as $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\})$, where ${\mathbf{z}}_{0}$ is the intermediate target feature (diffusion branch in Figure 2) and $\{{\mathbf{z}}_{i}\}$ are the input features corresponding to multi-view reference images (top two rows in Figure 2).
+We extract spatial features $\{{\mathbf{W}}_{i}\}$ from $\{{\mathbf{z}}_{i}\}$ using components of pre-trained U-Net itself, i.e., $F_{\text{standard}}({\mathbf{z}}_{i},{\mathbf{c}})$.
+To condition the diffusion branch on $\phi$, we learn a radiance field, denoted as FeatureNeRF, from $\{{\mathbf{W}}_{i},\pi_{i}\}$ in a feed-forward manner [102].
+The predicted FeatureNeRF is then rendered from the target viewpoint $\phi$ to obtain view-dependent feature map ${\mathbf{W}}_{y}$.
+In the main diffusion branch, we extract the intermediate feature map after the self and cross-attention layers, i.e., ${\mathbf{W}}_{{\mathbf{x}}}=g(s({\mathbf{z}}_{0}),\mathbf{c})$.
+We concatenate ${\mathbf{W}}_{{\mathbf{x}}}$ with the rendered features ${\mathbf{W}}_{y}$ and then project it into the original feature dimension using a linear layer.
+Thus, the pose conditioned transformer layer, $F_{\text{pose}}({\mathbf{z}}_{0},\{{\mathbf{z}}_{i},\pi_{i}\},\mathbf{c},\phi)$ performs:
+Figure 3: FeatureNeRF.
+We predict volumetric features $\overline{\mathbf{V}}$ for each 3D point in the grid using reference features $\{{\mathbf{W}}_{i}\}$ (Eqn.
+4).
+Given this feature, we predict the density $\sigma$ and color $rgb$ using a 2-layer MLP and use the predicted density $\sigma$ to render $\hat{\mathbf{V}}$ (which has been updated with text cross-attention $g$).
+The $rgb$ is only used to calculate reconstruction loss during training.
+${\mathbf{W}}_{i}=F_{\text{standard}}({\mathbf{z}}_{i},$ $\mathbf{c}),\hskip 5.69054pt{\mathbf{W}}_{y}=\text{FeatureNeRF}(\{{\mathbf{W}}_{i},\pi_{i}\},\mathbf{c},\phi)$ (3)
+$F_{\text{pose}}=f(l({\mathbf{W}}_{y}\oplus{\mathbf{W}}_{{\mathbf{x}}}))$
+where $l$ is a learnable weight matrix, which projects the feature into the input space of feed-forward layer $f$.
+We initialize $l$ such that the contribution from ${\mathbf{W}}_{y}$ is zero at the start of training.
 FeatureNeRF.
-Here, we describe the aggregation of individual 2D features ${\mathbf{W}}_{i}$ with 3D poses $\pi_{i}$ into a feature map ${\mathbf{W}}_{y}$ from pose $\phi$.
-Rather than learning NeRF in a feature space [97, 40], our focus is on learning 3D features that the 2D diffusion model can use.
-From target viewpoint $\phi$, for each point $\mathbf{p}$ on a target ray with direction $\mathbf{d}$, we project the point on the image plane of given views $\pi_{i}$ and denote projected locations as $\pi_{i}^{\mathbf{p}}$.
-We then sample from these coordinates on feature map ${\mathbf{W}}_{i}$, predict a feature for the 3D point, and aggregate them with function $\psi$:
-$\mathbf{V}_{i}=$ $\text{MLP}(\text{Sample}({\mathbf{W}}_{i};\pi_{i}^{\mathbf{p}}),\gamma(\mathbf{d}),\gamma(\mathbf{p})),\;{i=1,...,N}$ (4)
-$\bar{\mathbf{V}}=\psi(\mathbf{V}_{1},...,\mathbf{V}_{N}),$
-Figure 3: FeatureNeRF block.
-We predict volumetric features $\bar{\mathbf{V}}$ for each 3D point in the grid using reference features $\{W_{i}\}$ (Eqn.
-4).
-Given this feature, we predict the density $\sigma$ and color $rgb$ using a 2-layer MLP and use the predicted density to render $\hat{\mathbf{V}}$, which has been updated with text cross-attention $g$.
-The predicted $rgb$ is only used to calculate reconstruction loss during training.
+Here, we describe the aggregation of individual features ${\mathbf{W}}_{i}$ with poses $\pi_{i}$ into a feature map ${\mathbf{W}}_{y}$ from pose $\phi$.
+Given a target ray with direction $\mathbf{d}$ from target viewpoint $\phi$, we sample points $\mathbf{p}$ along the ray and project it to the image plane of each given view $\pi_{i}$.
+The projected coordinate is denoted as $\pi_{i}^{\mathbf{p}}$.
+We then sample the feature from this coordinate in ${\mathbf{W}}_{i}$, predict a feature for the 3D point $\mathbf{p}$, and aggregate the $N$ predicted features from each view with function $\psi$:
+$\mathbf{V}_{i}^{\mathbf{p}}=$ $\text{MLP}(\text{Sample}({\mathbf{W}}_{i};\pi_{i}^{\mathbf{p}}),\gamma(\mathbf{d}),\gamma(\mathbf{p})),\;{i=1,...,N}$ (4)
+$\overline{\mathbf{V}}^{\mathbf{p}}=\psi(\mathbf{V}_{1}^{\mathbf{p}},...,\mathbf{V}_{N}^{\mathbf{p}}),$
 where $\gamma$ is the frequency encoding.
-We use the weighted average [71] as the aggregation function $\psi$, where a linear layer predicts the weights based on $\mathbf{V}_{i}$, $\pi_{i}$, and target pose $\phi$.
-For each reference view, $\mathbf{d}$ and $\mathbf{p}$ are first transformed in the view coordinate space.
-We then predict the density and color of the point using a linear layer:
-$(\sigma,\mathbf{C})=\text{MLP}(\bar{\mathbf{V}}).$ (5)
-We also update the aggregated feature with text condition $\mathbf{c}$ using cross-attention:
-$\hat{\mathbf{V}}=\text{CrossAttn}(\bar{\mathbf{V}},\mathbf{c}),$ (6)
-and render this updated intermediate feature using the predicted density $\sigma$:
+We use the weighted average [71] as the aggregation function $\psi$, where a linear layer predicts the weights based on $\mathbf{V}_{i}$, $\pi_{i}$, and $\phi$.
+For each reference view, $\mathbf{d}$ and $\mathbf{p}$ are first transformed in the view coordinate space [102].
+Given the feature $\overline{\mathbf{V}}$ (superscript ${\mathbf{p}}$ is dropped for simplicity) for the 3D point, we predict the density and color using a linear layer:
+$(\sigma,\mathbf{C})=\text{MLP}(\overline{\mathbf{V}}),$ (5)
+and also update the aggregated feature with text prompt $\mathbf{c}$ using cross-attention:
+$\hat{\mathbf{V}}=\text{CrossAttn}(\overline{\mathbf{V}},\mathbf{c}).$ (6)
+We then render this updated feature volume using the predicted densities:
 ${\mathbf{W}}_{y}(r)=\sum_{j=1}^{N_{f}}T_{j}(1-\exp(-\sigma_{j}\delta_{j}))\hat{\mathbf{V}}_{j},$ (7)
-where $r$ is the target ray, $\hat{\mathbf{V}}$ is the aggregated feature of the point in the ray, $\sigma$ is the predicted density of that point, $N_{f}$ is the number of sampled points along the ray between the near and far plane of the camera, and $T_{j}=\exp(-\sum_{k=1}^{j-1}\sigma_{k}\delta_{k})$ handles occlusion until that point.
-Conditioning.
-To process the main target branch, we extract the intermediate 2D feature map after the self $s$ and cross-attention layers $g$, i.e., ${\mathbf{W}}_{{\mathbf{x}}}=g(s({\mathbf{z}}_{0}),\mathbf{c})$.
-We concatenate ${\mathbf{W}}_{{\mathbf{x}}}$ with the rendered features ${\mathbf{W}}_{y}$ and then project back into the original feature dimension using a linear layer.
-Thus, the output of the modified transformer layer is:
-$F_{\text{pose}}=f(l({\mathbf{W}}_{y}\oplus{\mathbf{W}}_{{\mathbf{x}}})),$ (8)
-where $l$ is a learnable weight matrix, which projects the feature into the space to be processed by the feedforward layer $f$.
-We initialize $l$ such that the contribution from ${\mathbf{W}}_{y}$ is zero at the start of training.
+where $r$ is the target ray, $\hat{\mathbf{V}}_{j}$ is the feature corresponding to the $j^{th}$ point along the ray, $\sigma_{j}$ is the predicted density of that point, $N_{f}$ is the number of sampled points along the ray between the near and far plane of the camera, and $T_{j}=\exp(-\sum_{k=1}^{j-1}\sigma_{k}\delta_{k})$ handles occlusion until that point.
+We build our FeatureNeRF design based on PixelNeRF [102] but update the aggregated features with text cross-attention and use learnable weighted averaging to aggregate reference view features.
+Through this layer, our focus is on learning 3D features that the 2D diffusion model can use rather than learning NeRF in a feature space [100, 42].
 Training loss.
@@
 $\mathcal{L}_{\text{rgb}}=\sum_{r}||M(r)(\mathbf{C}_{gt}(r)-\sum_{j=1}^{N_{f}}T_{j}(1-\exp(-\sigma_{j}\delta_{j}))\mathbf{C})||,$ (9)
-and two mask-based losses – (1) silhouette loss [70] $\mathcal{L}_{\text{s}}$ which forces the rendered opacity to be similar to object mask, and (2) background suppression loss [6, 7] $\mathcal{L}_{\text{bg}}$ which enforces the density of all background rays to be zero, as we only wish to model the object.
+and two mask-based losses as we only wish to model the object – (1) silhouette loss [70] $\mathcal{L}_{\text{s}}$ which forces the rendered opacity to be similar to object mask, and (2) background suppression loss [7, 8] $\mathcal{L}_{\text{bg}}$ which enforces the density of all background rays to be zero.
 $\mathcal{L}_{\text{s}}$ $=\sum_{r}||M(r)-\sum_{j=1}^{N_{f}}T_{j}(1-\exp(-\sigma_{j}\delta_{j}))||$ (10)
@@
 $\mathcal{L}=\mathcal{L}_{\text{diffusion}}+\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{bg}}\mathcal{L}_{\text{bg}}+\lambda_{\text{s}}\mathcal{L}_{\text{s}},$ (11)
-where $M$ is the object mask and $\lambda_{i}$ are hyperparameters to control the rendering quality of intermediate images vs the final denoised image.
-We keep $\lambda_{i}$ fixed across all experiments.
-We assume access to the object’s mask in the image, which is used to calculate the above losses.
-The three losses corresponding to FeatureNeRF are averaged across all pose-conditioned transformer layers.
+where $M$ is the object mask and $\lambda_{\text{rgb}}$, $\lambda_{\text{bg}}$, and $\lambda_{\text{s}}$ are hyperparameters to control the rendering quality of intermediate images vs.
+the final denoised image and are fixed across all experiments.
+We assume access to the object’s mask in the image, which is used to calculate these losses.
+The losses for FeatureNeRF are averaged across all pose-conditioned transformer layers.
 Figure 4: Qualitative comparison.
-Given a particular target pose, we show the qualitative comparison of our method with (1) Image editing methods SDEdit, InstructPix2Pix, and LEDITS++ which edit a NeRF rendered image from the input pose, (2) Vica-NeRF, a 3D editing method that trains a NeRF model for each input prompt, and (3) LoRA + Camera pose, our proposed baseline where we concatenate camera pose information to text embeddings during LoRA fine-tuning.
+Given a particular target pose, we show the qualitative comparison of our method with (1) Image editing methods SDEdit, InstructPix2Pix, and LEDITS++, which edit a NeRF-rendered image from the input pose, (2) ViCA-NeRF, a 3D editing method that trains a NeRF model for each input prompt, and (3) LoRA + Camera pose, our proposed baseline where we concatenate camera pose information to text embeddings during LoRA fine-tuning.
 Our method performs on par or better in keeping the target identity and poses while incorporating the new text prompt—e.g., putting a picnic table next to the SUV car ($1^{\text{st}}$ column)—and following multiple text conditions—e.g., turning the chair red and placing it in a white room ($3^{\text{rd}}$ column).
@@
 We show more sample comparisons in Figure 17 of Appendix.
-Figure 5: Qualitative samples with varying pose.
-Our method’s results with different text prompts and target poses as conditions.
-Our method learns the identity of custom objects while allowing the user to control the camera pose and text prompt for generating the object in new contexts, e.g., changing the background scene or object color and shape.
-In each row, the images were generated with the same seed while changing the camera pose around the object in a turntable manner.
+Figure 5: Qualitative samples with varying object viewpoint and text prompt.
+Our method learns the identity of custom objects while allowing the user to control the object viewpoint and generating the object in new contexts using the text prompt, e.g., changing the background scene or object color and shape.
+In each row, the images were generated with the same seed while changing the object viewpoint in a turntable manner.
+Note that each image in a row is independently generated.
 Figure 18 in the Appendix shows more such samples.
-Note that each image in a row is independently generated.
-We do not aim to generate multi-view consistent scenes.
 Inference.   During inference, to balance the text vs. reference view conditions in the final generated image, we combine text and image guidance [10] as shown below:
@@
 We also drop the text prompt with $10\%$ probability to be able to use classifier-free guidance.
-We provide more implementation details in Appendix C.
+We provide more implementation details in Appendix D.
 4 Experiments
 Dataset.
-For our experiments, we select concepts from the Common Objects in 3D (CO3Dv2) dataset [71], commonly used for novel view synthesis, and NAVI [33].
-Specifically, we select four categories with three instances from the CO3Dv2 dataset- car, chair, teddy bear, and motorcycle—as each instance is uniquely identifiable for these categories.
-From the NAVI dataset, we select two unique toy-like concepts.
-We use the camera pose provided in the dataset for the multi-view images.
-A representative image of each concept is shown in Figure 14, Appendix A.
-For each instance, we sample $\sim 100$ images and use half for training and half for evaluation.
+We select 14 custom objects from the CO3Dv2 [71] and NAVI [35] datasets.
+Specifically, we select $4$ categories with $3$ instances each from the CO3Dv2 dataset – car, chair, teddy bear, and motorcycle.
+From NAVI, we select $2$ unique, toy-like objects.
+A representative image of each concept is shown in the supplemental material.
+We use the camera poses provided in the dataset.
+For each instance, we sample $\sim 100$ images, using half for training and half for evaluation.
 The camera poses are normalized such that the mean of camera location is the origin, and the first camera is at unit norm [105].
@@
 Table 1: Human preference evaluation.
-Our method is preferred over all baselines for text alignment, image alignment to target concept, and photorealism except LoRA + Camera pose, which overfits the training images, as also shown in Figure 4.
+Our method is preferred over almost all baselines for text alignment, image alignment to the target concept, and photorealism.
+We find that LoRA + Camera pose overfits the training images, as shown in Figure 4.
 Method Angular error Camera center error
@@
 LoRA + Camera pose 41.14 0.305
-Table 2: Camera pose accuracy in generated images by ours and the LoRA + Camera pose baseline method.
-We observe that the baseline usually overfits to training images and does not respect the target pose with new text prompts.
+Table 2: Accuracy of object viewpoint condition in generated images by ours and the LoRA + Camera pose baseline method.
+We observe that the baseline usually overfits the training images and does not respect the target viewpoint condition with new text prompts.
 Baselines.
-While no prior method targets our exact task, we use three types of related baselines – (1) 2D image editing methods, which aim to preserve details of the input image and thus keep the object in the same pose as the input image.
-This includes three recent and publicly available methods: LEDITS++ [8], InstructPix2Pix [9], and SDEdit [52] with Stable Diffusion -1.5 (and SDXL in Appendix A).
-As image editing methods by themselves do not support camera viewpoint manipulation, we first render a NeRF model [87] in the target pose and then edit the rendered image.
-(2) Customization-based method, LoRA+Camera pose, where we modify LoRA [76, 32] by concatenating the camera pose information to the text embeddings, following recent work Zero-1-to-3 [49].
-(3) VICA-NeRF [20], a 3D editing method that trains a NeRF for each new text prompt.
-In Appendix C, we provide more details on implementation and hyperparameters for each baseline.
+We compare with three types of relevant baselines – (1) First, we compare against 2D image editing using $3$ recent, publicly available methods – LEDITS++ [9], InstructPix2Pix [10], and SDEdit [53] with Stable Diffusion-1.5 (and SDXL in Appendix 4.1).
+As image editing methods do not inherently support viewpoint manipulation, we first render a NeRF [87] of the input scene with the target viewpoint and then edit the rendered image.
+(2) Secondly, we use a customization-based method, LoRA+Camera pose, where we modify LoRA [76, 34] by concatenating the camera pose to the text embeddings, following recent work Zero-1-to-3 [50].
+(3) Lastly, we test ViCA-NeRF [22], a 3D editing method that trains a NeRF for each new text prompt.
+In Appendix D, we provide more details on implementation and hyperparameters for each baseline.
 Evaluation metrics.
@@
 We then manually inspect them to remove implausible or overly complicated text prompts [93].
-Table 5 in Appendix B lists all the evaluation prompts.
-For a quantitative comparison, we primarily use a pairwise human preference study.
-We compare our method against each baseline to measure image alignment to target concept, alignment to input text prompt, and photorealism of generated images.
-In total, we collect $\sim 1000$ responses per pairwise study using Amazon Mechanical Turk.
-We also show the performance of our method and baselines on other standard metrics like CLIP Score [67] and DINOv2 [59] image similarity [74] to measure the text- and image-alignment.
-To measure whether the object in generated images corresponds to the input camera pose for our method and the LoRA + Camera pose baseline, we use a pretrained model, RayDiffusion [101], to predict the poses from generated images and calculate its error relative to the ground truth camera poses.
-More details about evaluation are provided in Appendix B.
+Table 5 in Appendix C lists all the evaluation prompts.
+We evaluate (1) the customization quality of the generated image and (2) its adherence to the specific pose.
+First, to measure customization quality, we use a pairwise human preference study.
+A successful customization is comprised of several aspects: alignment to the target concept, alignment to the input text prompt, and photorealism of the generated images.
+In total, we collect $\sim 1000$ responses per pairwise study against each baseline using Amazon Mechanical Turk.
+We also evaluate our method and baselines on other standard metrics like CLIP Score [67] and DINOv2 [61] image similarity [74] to measure the text and image alignment.
+To measure whether the generated custom object adheres to the specified viewpoint, we use a pre-trained model, RayDiffusion [106], to predict the pose from the generated images and calculate its error relative to the input camera pose.
+More details about evaluation are provided in Appendix C.
 Figure 6: Quantitative comparison.
@@
 4.1 Results
-Figure 7: Extrapolating camera pose from training views.
-Our method can generalize to different camera poses, including viewpoints not within the training distribution.
+Figure 7: Extrapolating object viewpoint from training viewpoints.
+Our method can generalize to different viewpoints, including those not within the training distribution.
 Top left: We vary the focal length from ×$\times 0.8$ to ×$\times 1.4$ of the original focal length.
@@
 Generation quality and adherence.
-First, we measure the quality of the generation – adherence to the text prompt, the identity preservation to the customized objects, and photorealism – irrespective of the camera pose.
-For the comparison, we generate $18$ images per prompt on $6$ target camera poses, totaling $288$ images per concept.
+First, we measure the quality of the generation – adherence to the text prompt, the identity preservation to the customized objects, and photorealism – irrespective of the object viewpoint.
+Recall that for each concept, we curate 16 prompts.
+For each prompt, we generate $3$ images at each viewpoint, covering $6$ target viewpoints, resulting in $288$ images per concept.
 Table 1 shows the pairwise human preference for our method vs.
 baselines.
-Our method is preferred over all baselines except LoRA + Camera pose, which we observe to overfit on training images.
+Our method is preferred over all baselines except LoRA + Camera pose, which we observe to overfit on training images, thus producing higher image alignment.
 Figure 6 shows the CLIP vs.
-DINO scores for all methods and object categories.
-Ideally, a method should have both a high CLIP score and a DINO score, but often, there is a trade-off between text- and image alignment.
-Our method has on-par or better text alignment relative to the baselines while having better image alignment.
+DINO scores for all methods.
+Ideally, a method should have both a high CLIP score and a DINO score, but often, there is a trade-off between text and image alignment.
+Our method has on-par or better text alignment relative to the baselines, while having better image alignment.
 We observe that image-editing baselines often require careful hyperparameter tuning for each image.
 We select the best-performing hyperparameters and keep them fixed across all experiments.
-We use the $\sim 50$ validation camera poses not used during training for evaluation and randomly perturb the camera position or focal length.
-Figure 11 in Appendix B shows sample training and perturbed validation camera poses for the car object.
-Camera pose accuracy.
-Previously, we have measured our method purely on image customization benchmarks.
-Next, we evaluate the camera pose accuracy as well.
-Table 2 shows the camera pose accuracy of the generated images in terms of mean angular error and camera center error.
-We observe that LoRA + Camera pose baseline overfits on training images and can fail at generating images in the correct pose with new text prompts during inference.
-We evaluate this on validation camera poses of concepts from the CO3Dv2 dataset with the camera’s principal axis pointing towards the object at the scene’s center.
-This is because RayDiffusion has been trained on this setup of the CO3Dv2 dataset and fails on unique objects.
+The camera pose corresponding to the target object viewpoint is uniformly sampled from $\sim 50$ validation poses not used during training.
+We also randomly perturb the camera position or focal length.
+Figure 13 in Appendix C shows sample training and perturbed validation camera poses for the car object.
+Accuracy of object viewpoint.
+Previously, we evaluated our method purely on image customization benchmarks.
+Next, we evaluate the accuracy of the object viewpoint conditioning.
+Table 2 shows the mean angular error and camera center error between the generated object’s pose, predicted using RayDiffusion [106], and the input pose.
+We only compare with LoRA + Camera pose, as only this baseline takes the camera pose for the target object viewpoint as input.
+We observe that it often overfits training images and fails to generate the object in the correct viewpoint with new text prompts.
+We evaluate this metric only on the objects from the CO3Dv2 dataset with validation camera poses, as RayDiffusion has been trained on CO3Dv2 and struggles with other unique objects.
 Qualitative comparison.
 We show the qualitative comparison of our method with the baselines in Figure 4.
-As we can see, image-editing-based methods often fail at generating photorealistic results.
-In the case of LoRA + Camera Pose, we observe that it fails to generalize and overfits to the training views ($5^{\text{th}}$ row Figure 4).
-Finally, the 3D editing-based method Vica-NeRF maintains 3D consistency but generates blurred images for text prompts that change the background scene.
-Figure 5 shows samples with different text prompts and target camera poses for our method.
-Generalization to novel camera poses.
-Since our method learns a 3D radiance field, we can also extrapolate to unseen camera poses at inference time as shown in Figure 7.
+We observe that image editing methods can fail to generate photorealistic results.
+In the case of LoRA + Camera pose, it fails to generalize and overfits to the training views ($5^{\text{th}}$ row).
+Finally, the 3D editing-based method ViCA-NeRF maintains 3D consistency but generates blurred images, especially for text prompts that change the background.
+Figure 5 shows more samples with different text prompts and object viewpoints for our method.
+Additional comparison to customization + 3D-aware image editing.
+We further compare against a two-stage approach that first generates an image of the custom object using LoRA+DreamBooth [76, 74] and then edits the object to a target viewpoint using two recent 3D-aware image editing methods, Image Sculpting [101] and Object3DIT [55].
+For each prompt, we generate $3$ images, then edit and rotate the object to 6 different viewpoints.
+This results in 288 images per concept, similar to our evaluation setting.
+We compare against this on only the three car objects since Image-Sculpting uses Adobe Photoshop’s generative fill [1] as one of the intermediate steps, which requires manually inpainting each image.
+The CLIP scores for Image Sculpting and Object3DIT are $0.26$ and $0.27$, respectively, compared to our score of $0.25$.
+However, their DINO scores at $0.24$ and $0.40$ are substantially lower than our $0.48$.
+As shown in Figure 8, both methods lead to lower-fidelity results.
+Object3DIT struggles in many scenarios due to its training on a synthetic dataset, and Image Sculpting’s performance is highly dependent on single image-to-3D methods like Zero-1-to-3 [50] used in its pipeline.
+Generalization to novel viewpoints.
+Since our method learns a 3D radiance field, we can also extrapolate to unseen object viewpoints at inference time as shown in Figure 7.
 We generate images while varying the camera distance from the object (scale), focal length, or camera position along the horizontal and vertical axis.
+Figure 8: Comparison to 3D-aware editing methods.
+We first generate the image of the custom object using LoRA+DreamBooth (shown as an inset) and then use the 3D-aware editing method to edit and rotate the object to a target viewpoint.
+We show qualitative samples generated by our method ($3^{\text{rd}}$ column) with approximately the same target viewpoint as input.
+Object3DIT and Image-Sculpting lead to lower fidelity edits than images generated by our method ($3^{\text{rd}}$ column) with the target viewpoint directly as the input condition.
 Figure 9: Applications.
 $1^{\text{st}}$ row: Our method can be combined with other image editing methods as well.
-We use SDEdit with our method to in-paint the rubber duck in different poses while keeping the same background.
-$2^{\text{nd}}$ row: We can generate interesting panorama shots by controlling the camera pose of the object in each grid independently.
+We use SDEdit with our method to in-paint the car and rubber duck from different viewpoints while keeping the same background.
+$2^{\text{nd}}$ row: We can generate interesting panorama shots by controlling the object viewpoint independently in each grid.
 $3^{\text{rd}}$ row: We can also compose the radiance field predicted by FeatureNeRF to control the relative pose while generating multiple instances of the object.
@@
 Our method can be combined with existing image editing methods as well.
-Figure 8a shows an example where we use SDEdit [52] to in-paint the object in varying poses while keeping the same background.
-We can also generate interesting panoramas using MultiDiffusion [3], where the object’s placement in each grid is controlled by our method, as shown in Figure 8b.
-Moreover, since we learn a 3D consistent FeatureNeRF for the new concept, we can compose multiple instances of the object in feature space [84], with each instance in a different camera pose.
-Figure 8c shows an example of two teddy bears facing each other sitting on the armchair.
+Figure 9a shows an example where we use SDEdit [53] to in-paint the object with different viewpoints while keeping the same background.
+We can also generate interesting panoramas using MultiDiffusion [4], where the object viewpoint in each grid is controlled by our method, as shown in Figure 9b.
+Moreover, since we learn a 3D consistent FeatureNeRF for the new object, we can compose multiple instances of the object [84], with each instance in a different viewpoint.
+Figure 9c shows an example of two teddy bears facing each other and sitting on armchairs.
 Here, we additionally use DenseDiffusion [44] to modulate the attention maps and guide the generation of each object instance to only appear in the corresponding region predicted by FeatureNeRF.
@@
 Not having mask-based losses (Eqn.
-11) leads to overfitting on training images and decreases the text alignment.
+10) leads to overfitting on training images and decreases the text alignment.
 The worst performing metrics are grayed.
@@
 10 from training, we observe a decrease in text alignment and overfitting on training images as shown in Table 3.
-Figure 12 in Appendix A shows qualitatively that the model generates images with backgrounds more similar to the training views.
+Figure 14 in Appendix 4.1 shows qualitatively that the model generates images with backgrounds more similar to the training views.
 This is also reflected by the higher similarity between generated images and background regions of the training images ($3^{\text{rd}}$ column Table 3) compared to our final method.
@@
 Figure 10: Limitations.
-Our method can fail when extrapolating camera poses far from the training image camera poses, e.g., changing focal length (top left) or translating camera s.t.
-the object is not in the center (top right) as the pre-trained model is often biased towards generating the object in the center.
-Also, it can fail to follow the input text prompt or the exact camera pose when multiple objects are composed in a scene (bottom row).
+Our method can occasionally fail when the target object viewpoint deviates far from the training images, e.g., reducing the focal length too much (top left) or rendering the object off-center (top right), as the pre-trained model is often biased towards generating the object in the center.
+Also, it can fail to follow the input text prompt or the exact object viewpoint when multiple objects are composed in a scene (bottom row).
+We show more results and ablation experiments in the Appendix, including performance with predicted camera viewpoints.
 5 Discussion and Limitations
-We introduce a new task of customizing text-to-image models with camera viewpoint control.
-Our method jointly learns 3D feature prediction modules and adapts 2D diffusion attention modules to be conditioned on these features.
-This enables synthesizing the object with new text prompts and precise object pose control.
-While our method outperforms existing image editing and model customization approaches, it still has several limitations, as we discuss below.
+We introduce a new task of customizing text-to-image models with object viewpoint control.
+Our method learns view-dependent object features in the intermediate feature space of the diffusion model and conditions the generation on them.
+This enables synthesizing the object with varying object viewpoints while controlling other aspects through text prompts.
 Limitations.
-As we show in Figure 9, our method struggles at generalizing to extreme camera poses that were not seen during training and resorts to either changing the object identity or generating the object in a seen pose.
-We expect this to improve by adding more camera pose variations during training.
-An additional case when our method struggles to follow the camera pose is when the text prompt adds multiple new objects.
-We hypothesize that in such challenging scenarios, the model is biased towards generating front views, often seen in training.
-Also, our proposed pose-conditioning module is trained with a finetuning-based method.
-Exploring pose-conditioning in a zero-shot customization and editing methods [16, 24] may help reduce the time and computation as well.
-Here, we focus on enabling camera view control while generating rigid objects.
+Though our method outperforms existing image editing and model customization approaches, it still has several limitations.
+As we show in Figure 10, our method occasionally struggles at generalizing to extreme viewpoints that were not seen during training and resorts to either changing the object identity or generating the object in a seen viewpoint.
+We expect this to improve by adding more viewpoint variations during training.
+Our method also sometimes struggles to follow the input viewpoint condition when the text prompt adds multiple objects to the scene.
+We hypothesize that in such challenging scenarios, the model is biased towards generating object-centric front views, as seen in its original training data.
+Also, we fine-tune the model for each custom object, which takes computation time ($\sim$ 40 minutes).
+Exploring pose-conditioning in a zero-shot, feed-forward manner [18, 26] may help reduce the time and computation.
+Finally, we focus on enabling viewpoint control for rigid objects.
 Future work includes extending this conditioning to handle dynamic objects that change the pose in between reference views.
@@
 Appendix
-In Appendix A, we show more results and baseline comparisons.
-In Appendix B, we provide details regarding the evaluation and human preference study.
-Finally, in Appendix C, we describe all the implementation details of our method and baselines.
-Appendix A Results
-More baselines.
-We add a comparison to SDEdit with the Stable Diffusion-XL model here, which we found to perform worse or on par with the Stable Diffusion-1.5 version of the model.
-Figure 10 shows the quantitative comparison.
-Per category comparison.
-We show text-alignment and image-alignment scores of our method and baselines for each category in Figure 10 while varying the text guidance scale for each method from $5.0$ to $10.0$, except for LEDITS++ [8] which recommends varying the concept editing guidance scale from $10.0$ to $15.0$.
-For each guidance scale, we generate $288$ images.
-Our method usually results in a higher CLIP score than the baselines, indicating better text alignment.
-For all categories, we show the linear fit curve over the different guidance scales, and our method lies at the Pareto-frontal compared to the baseline methods, either performing similarly or better.
+In Appendix A and Section B, we show more ablation and results.
+In Appendix C, we provide details regarding the evaluation and human preference study.
+Finally, in Appendix D, we describe all the implementation details of our method and baselines.
+Appendix A Ablation
 Using predicted masks for background losses.
 In our main paper, we calculated the background losses (Eqn.
-11) using ground truth masks from the dataset during training.
+10) using ground truth masks from the dataset during training.
 We show in Table 4 that replacing these with predicted masks results in similar performance.
@@
 Example generations and their comparison to our main method are shown in Figure 15.
-Qualitative comparison.   We show more qualitative comparisons between our method and baselines in Figure 15. Figure 16 shows more samples from varying camera poses for our method.
+Using predicted camera viewpoints.
+Here, we use COLMAP to predict the camera viewpoints using only the $50$ training images.
+Though it requires longer training, $2000$ steps compared to $1600$ for our method, the final performance is comparable to our original method.
+The CLIP and DINO scores remain similar while the mean angular error increases slightly to $17.12$ compared to $16.14$ of our final method on $10$ objects (COLMAP fails to run on $2$ car objects).
+Qualitative samples and their comparison to our final method are shown in Figure 11.
+Appendix B More Results
+Per category comparison.
+We show text-alignment and image-alignment scores of our method and baselines for each category in Figure 12 while varying the text guidance scale for each method from $5.0$ to $10.0$, except for LEDITS++ [9] which recommends varying the concept editing guidance scale from $10.0$ to $15.0$.
+For each guidance scale, we generate $288$ images.
+Our method usually results in a higher CLIP score than the baselines, indicating better text alignment.
+For all categories, we show the linear fit curve over the different guidance scales, and our method lies at the Pareto-frontal compared to the baseline methods, either performing similarly or better.
+Comparison with SDEdit+SDXL.
+We add a comparison to SDEdit with the Stable Diffusion-XL model here, which we found to perform worse or on par with the Stable Diffusion-1.5 version of the model.
+Figure 12 shows the quantitative comparison.
+Comparison to IP-Adapter + ControlNet.
+Here, we compare against IP-Adapter [99] + depth ControlNet [107] for object viewpoint control.
+Given a target viewpoint, we render the image using a trained NeRF model of the object and use its predicted depth as a condition in depth ControlNet.
+We select the recommended hyperparameters that also worked best qualitatively, i.e., $0.6$ scale for the IP-Adapter and $0.5$ as ControlNet conditioning scale.
+The CLIP and DINO scores for the baseline are $0.251$ and $0.497$, respectively, similar to our method’s CLIP and DINO scores of $0.253$ and $0.481$.
+But in pairwise user-study comparison, our method is preferred by $60\pm 3\%$ and $75.67\pm 2.6\%$ for text-alignment and image-alignment.
+Qualitatively, we also observe that applying multiple conditionings, including depth and image, leads to lower text alignment as shown in Figure 17 ($5^{\text{th}}$ row).
 Comparison to LoRA.
@@
 In comparison, our CLIP and DINO scores are $0.253$ and $0.481$, respectively.
-Though the CLIP score is marginally lower, our method performs better in preserving the target concept while allowing additional camera pose control for the custom object.
+Though the CLIP score is marginally lower, our method performs better in preserving the target concept while allowing additional object viewpoint control for the custom object.
+Qualitative comparison.   We show more qualitative comparisons between our method and baselines in Figure 17. Figure 18 shows more samples from varying camera poses for our method.
 Method Text Alignment Image Alignment Camera-pose Accuracy
@@
 When varying the number of views, the performance gradually drops w.r.t.
-camera pose accuracy and image alignment.
-Appendix B Evaluation.
+camera-pose accuracy and image alignment.
+Appendix C Evaluation.
 Evaluation text prompts.
@@
 We crop the object region in the training images using masks to measure the similarity only with the target concept.
-Appendix C Implementation Details
-C.1 Our Method
+Figure 11: Training with predicted camera viewpoints. We observe similar performance with predicted cameras using $50$ training images.
+Appendix D Implementation Details
+D.1 Our Method
 We fine-tune a pretrained Stable Diffusion-XL $32$-bit floating point model with a batch size of $4$.
@@
 During inference, we use an image guidance scale of $3.5$ and a text guidance scale of $7.5$ in Eqn.
-13.
+12.
 All images for evaluation are generated with $50$ sampling steps using the default Euler scheduler [37].
@@
 We observe that the model has a bias towards generating front-facing objects.
-C.2 Baselines
+D.2 Baselines
 For 2D image editing methods, we first render the image given the target camera pose using a trained NeRF model [87].
@@
 We train the model for a total of $2000$ iterations.
+Appendix E Change log
+v1:   Original draft.
+v2:   Aditional comparison to 3D-editing based methods in Section 4.1 and more ablation experiments in Appendix A. Updated citations and writing.
 Object Category Evaluation Prompt